### v1

# Hardware-Compatible Single-Shot Feasible-Space Heuristics for Solving the Quadratic Assignment Problem

Abstract
Research into the development of special-purpose computing architectures designed to solve quadratic unconstrained binary optimization (QUBO) problems has flourished in recent years. It has been demonstrated in the literature that such special-purpose solvers can outperform traditional CMOS architectures by orders of magnitude with respect to timing metrics on synthetic problems. However, they face challenges with constrained problems such as the quadratic assignment problem (QAP), where mapping to binary formulations such as QUBO introduces overhead and limits parallelism. In-memory computing (IMC) devices, such as memristor-based analog Ising machines, offer significant speedups and efficiency gains over traditional CPU-based solvers, particularly for solving combinatorial optimization problems. In this work, we present a novel local search heuristic designed for IMC hardware to tackle the QAP. Our approach enables massive parallelism that allows for computing of full neighbourhoods simultaneously to make update decisions. We ensure binary solutions remain feasible by selecting local moves that lead to neighbouring feasible solutions, leveraging feasible-space search heuristics and the underlying structure of a given problem. Our approach is compatible with both digital computers and analog hardware. We demonstrate its effectiveness in CPU implementations by comparing it with state-of-the-art heuristics for solving the QAP.
1 Introduction
The quadratic assignment problem (QAP) was originally introduced in 1957 by Koopmans and Beckmann [17]. In the QAP, the goal is to assign a number of facilities to an equal number of locations such that the cost associated with flows (e.g., energy or goods) between the facilities is minimized. This problem has remained a popular research topic in the literature due to its ubiquity in various domains, such as facility location, scheduling, process communications, warehouse management, and other fields; see, for example, Ref. [11, 18, 7, 5]. In its native formulation, the solution space of the QAP is represented as comprising all possible permutations, where each permutation represents distinct one-to-one assignments of facilities to locations. An instance of the QAP may be denoted as $\mathrm{QAP}(F,D)$, where $F=[F_{ij}]_{n\times n}$ and $D=[D_{kl}]_{n\times n}$ represent the flow and the distance matrices between $n$ facilities and $n$ locations, respectively. The goal is to find the assignment of facilities to locations such that the total flow cost is minimized. An assignment decision is feasible if and only if it is a permutation. Given a set $\Pi_{n}$ of all permutations of $n$ facilities, the aim of $\mathrm{QAP}(F,D)$ is to find an optimal permutation $\pi^{*}\in\Pi_{n}$ such that
$\pi^{*}\in\operatorname*{Argmin}_{\begin{subarray}{c}\pi\in\Pi_{n}\end{subarray}}\left\{\sum_{i=1}^{n}\sum_{j=1}^{n}F_{ij}D_{\pi(i)\pi(j)}\right\}.$ (1.1)
Research on combinatorial optimization has focused primarily on benchmarking heuristics and algorithms running on traditional CMOS hardware: mainly CPUs, but increasingly also GPUs, field-programmable gate arrays (FPGA), and even some custom application-specific integrated chip (ASIC) implementations of various algorithms. The performance of CPUs has doubled approximately every two years since the early 1970s, an empirical observation that has become known as Moore’s law. This steady exponential increase in available computing power is expected to slow down in the current decade because the density of transistors in integrated circuits cannot grow indefinitely beyond fundamental physical limits at the nanometer scale. Relatedly, Dennard’s scaling law states that, as transistors shrink, their power density remains constant, allowing for increased computing efficiency per chip area. Combined with Moore’s law, it has resulted in a doubling in performance per joule every 18 months. However, Dennard’s scaling law began to break down in the mid-2010s. For this reason, attention is shifting toward alternative computing paradigms, such as GPUs, FPGAs, and ASICs, as well as to devices that exploit quantum mechanical effects, and devices that exploit analog effects to perform very energy-efficient computations [6, 10]. D-Wave Systems’ quantum annealer [16], Fujitsu’s Digital Annealer [21, 1], NTT’s coherent Ising machine [15], Hitachi’s FPGA-based Ising computers [25, 33, 24], and Toshiba’s simulated bifurcation machine [14] are all examples of new technologies that are expected to benefit the field of computationally intensive discrete optimization problems, and can be generally referred to as in-memory computing (IMC) hardware accelerators. Application-driven analog IMC hardware devices that embed problem data within a crossbar structure are another promising approach [29, 4]. In-memory computing hardware accelerators are known for their energy-efficient computation [19, 26, 34] and, when combined with application-specific algorithms, can fully exploit parallelism, leading to significantly improved computation times and reduced computational energy consumption. In-memory computing hardware typically benefits heuristics that can tolerate limited precision, but has limitations, for example, implementing algorithms with complex decision heuristics, such as backtracking.
The majority of the novel optimization hardware solutions that have been developed in recent years aim to solve instances of quadratic unconstrained binary optimization (QUBO) problems, or using the equivalent Ising formulation. Such devices are promoted as fast general-purpose optimization solvers because a wide variety of hard problems can be mapped to QUBO form in polynomial time [20]. The literature that has showcased the progress of these new devices has not, however, demonstrated their efficiency in handling various practical industrial optimization problems. New computational devices are still maturing and the technological limitations, such as limited numbers of bits, qubits, connectivity, low precision, and noise will likely be addressed in the future. Mapping real-world problems into QUBO form presents other challenges that cannot be fixed by hardware advances. The reformulation into QUBO, particularly for constrained problems, often results in significantly harder problems by creating a much larger, more rugged solution landscape that includes both feasible and infeasible solutions [8]. These infeasible solutions are typically cast as a form of penalty in the objective function [13], which results in an objective function with many local optima. Searching in a mixed landscape of feasible and infeasible solutions has the key disadvantage that intermediary infeasible solutions must be traversed in order to move from one feasible solution to another, and even sophisticated algorithms struggle to find good solutions [22]. Furthermore, these devices face challenges with constrained problems like the QAP, where mapping to binary formulations such as QUBO introduces overhead and limits parallelism. Recent work has shown that state-of-the-art heuristics for solving the QAP can be mapped to FPGA and custom ASIC architectures, making use of massive parallelization to evaluate all possible local updates and guide a stochastic local search heuristic [2, 19].
In this paper, we present a novel local search heuristic, designed for IMC hardware, that is designed to explore the feasible space in the binary formulation of the QAP, and perform a single-shot evaluation of the full solution neighbourhood. Our approach enables massive parallelism by enabling a one-step, single-shot evaluation of the full neighbourhood or local gradient, enabling optimal local decisions, which can be made efficiently on a single computing core. When parallel search is discussed in the context of QAP heuristics, it is typically associated with leveraging multiple parallel processors to run independent heuristic searches [23]. In our approach, the full gradient is evaluated in parallel within a single iteration of the algorithm; see, for example, Ref. [26]. Binary solutions remain feasible by selecting local moves that lead to neighbouring feasible solutions, enabling the use of feasible-space search heuristics. As a result, our approach achieves an exponential reduction in search space compared with QUBO implementations, along with a much simpler energy landscape. Additionally, our method is compatible with both digital computers and IMC hardware accelerators. We demonstrate its effectiveness in CPU implementations by comparing it with various heuristics for solving the QAP.
1.1 Notation
[…]
1.2 Contributions
The contributions of this paper are as follows.
  1. 1.
    We provide an efficient construction of the full neighbourhood at a given feasible binary solution.
  2. 2.
    We design a parallel evaluation framework that computes objective differences across all neighbouring solutions in one iteration.
  3. 3.
    We introduce a framework tailored to the architectural requirements of application-driven IMC hardware.
1.3 Outline
Section 2 provides a background on the QAP and a discussion on the motivation for using the feasible-space method. In Section 3, we show formulas that leverage problems’ underlying structure to optimize performance. Section 4 introduces the novel full-neighbourhood search that explores an entire neighbourhood in parallel, potentially benefiting from IMC hardware. The benchmarking heuristics that we developed and tested, along with our experimental results, are outlined in Section 5. We summarize our work in Section 6.
2 Background
In this section, we review existing approaches for solving the QAP and explain the rationale behind pursuing the feasible-space method. Being an NP-hard problem, the QAP is commonly tackled using local-search heuristic algorithms, such as those presented in Refs. [9, 3, 12]. They are CPU-based solvers that offer fast runtimes, and have been shown to produce near-optimal solutions. Recently, hardware accelerators have been proposed as a means to solve challenging combinatorial optimization problems; see, for example, Ref. [8]. Problems are often formulated in the QUBO form in order to use these devices. Some examples of problems that have been solved by such devices using a QUBO formulation include the max-cut problem, graph colouring problem [20], $k$-satisfiability problem [30, 4], and the travelling salesperson problem [32]. It also possible to map the QAP to a binary formulation, which has been used to solve it with QUBO-based IMC hardware accelerators. In this section, we provide an overview of the binary formulation of the QAP, and discuss the motivation for designing the feasible-space method within the binary space.
2.1 Binary Formulation of the QAP
To solve the QAP using an Ising hardware solver to solve a QUBO problem, it needs to be cast into a binary formulation. Let $x_{ik}$ be a binary decision variable that is equal to $1$ if facility $i$ is assigned to location $k$, and 00 otherwise. The QUBO formulation of the problem (1.1) can be written using the following objective function $E$:
$\min_{x\in\{0,1\}^{n^{2}}}\sum_{i=1}^{n}\sum_{k=1}^{n}\sum_{j=1}^{n}\sum_{l=1}^{n}D_{ij}F_{kl}x_{ik}x_{jl}+\lambda\left[\sum_{k=1}^{n}\left(\sum_{i=1}^{n}x_{ik}-1\right)^{2}+\sum_{i=1}^{n}\left(\sum_{k=1}^{n}x_{ik}-1\right)^{2}\right].$ (2.1)
A penalty multiplier $\lambda$ is needed in order to ensure that only one facility is assigned to each location and vice versa. The parameter $\lambda$ should be chosen sufficiently large such that for every infeasible assignment $x_{\mathrm{inf}}$, there is at least one feasible assignment $x_{\mathrm{f}}$, where $E(x_{\mathrm{inf}})>E(x_{\mathrm{f}})$. This yields $E(x_{\mathrm{inf}})>\lambda$ because the flow and the distance matrices are non-negative and the assignment feasibility violation value is at least two. Further, let $u$ be an upper bound such that $E(x_{\mathrm{f}})<u$ for every feasible solution (i.e., permutation) $x_{\mathrm{f}}$. Setting $\lambda>u$ would ensure every non-permutation solution evaluates to a value greater than the upper bound $u$, and, as a result, the objective values of all permutation solutions are smaller than all non-permutation solutions’ objective values. The parameter $u$ is determined by sending the largest outgoing and incoming flows of each facility over the longest distances. This results in a loose bound on $\lambda$, and it might be possible to find a tighter bound. Thus, the penalty multiplier can be set to $\lambda=\alpha u$, and the parameter $\alpha\in[0,1]$ is tuned for optimal performance of the algorithm. Of note, with a tight bound on $\lambda$, all feasible solutions will be local optima in the extended binary solution space that also considers infeasible solutions.
2.1.1 Solution Space
There are multiple issues that contribute to making the binary formulation of the QAP much more challenging to solve than the native formulation (1.1), at least when using a standard single bit-flip QUBO solver. One of the main issues with the binary formulation is the exponential increase of the solution space. In the native QAP formulation, the solution space is given by the set of all permutations, $\Pi_{n}$. For a given value of $n$, there are $n!$ possible solutions, whereas in the binary formulation, the solutions are expressed as binary vectors of size $n^{2}$; therefore, the total solution space is has a size of $2^{n^{2}}$. Table 2.1 shows a comparison of the size of the solution space between the native QAP and the binary formulations for several problem sizes.
Number of Facilities
and Locations ($n$)
Number of
Binary Variables ($n^{2}$)
Solution Space:
Native ($n!$)
Solution Space:
Binary ($2^{n^{2}}$)
20 400 $\sim 10^{18}$ $\sim 10^{120}$
40 1600 $\sim 10^{48}$ $\sim 10^{482}$
60 3600 $\sim 10^{82}$ $\sim 10^{1084}$
80 6400 $\sim 10^{119}$ $\sim 10^{1927}$
100 10,000 $\sim 10^{158}$ $\sim 10^{3010}$
Table 2.1: Number of variables and size of the solution space in the native and binary formulations for various problem sizes.
The exponentially larger solution space of the QUBO formulation comprises mostly infeasible solutions, for example, solutions where a single location has two separate facilities assigned to it. In standard QUBO local-search heuristics, solutions are updated through single bit-flips, and since there are $n^{2}$ binary variables, there are exactly $n^{2}$ configurations accessible from any given solution via single bit-flips. Given a feasible starting configuration, none of the $n^{2}$ single bit-flip moves lead to another feasible configuration; in fact, they result in a solution penalized by the term $\lambda\left[\sum_{k=1}^{n}\left(\sum_{i=1}^{n}x_{ik}-1\right)^{2}+\sum_{i=1}^{n}\left(\sum_{k=1}^{n}x_{ik}-1\right)^{2}\right]$ in the QUBO objective function (2.1). It can easily be shown that to move via single bit-flips between feasible solutions in the binary space, it is necessary to visit at least three infeasible solutions before reaching another feasible solution. A feasible binary solution can be represented as an $n\times n$ permutation matrix by rearranging the $n^{2}$ solution vector, and a local move to another feasible solution corresponds to swapping two columns of the permutation matrix. The minimal way to do this is via four individual bit flips.
In the native space of QAP problems of size $n=3$, the basic move involves simply swapping two elements of a solution as follows:
$\begin{pmatrix}{\color[rgb]{1,0,0}3}\\ {\color[rgb]{1,.5,0}1}\\ {\color[rgb]{0,0,1}2}\end{pmatrix}\quad\rightarrow\quad\begin{pmatrix}{\color[rgb]{1,.5,0}1}\\ {\color[rgb]{1,0,0}3}\\ {\color[rgb]{0,0,1}2}\end{pmatrix}$
In single bit-flip QUBO heuristics, the equivalent move between two feasible QAP solutions must traverse at least three infeasible solutions (infeasible solutions are shown in grey and black):
$\begin{pmatrix}{\color[rgb]{1,0,0}0}\\ {\color[rgb]{1,0,0}0}\\ {\color[rgb]{1,0,0}1}\\ {\color[rgb]{1,.5,0}1}\\ {\color[rgb]{1,.5,0}0}\\ {\color[rgb]{1,.5,0}0}\\ {\color[rgb]{0,0,1}0}\\ {\color[rgb]{0,0,1}1}\\ {\color[rgb]{0,0,1}0}\end{pmatrix}\quad\rightarrow\quad\begin{pmatrix}{\color[rgb]{0,0,0}1}\\ {\color[rgb]{.75,.75,.75}0}\\ {\color[rgb]{.75,.75,.75}1}\\ {\color[rgb]{.75,.75,.75}1}\\ {\color[rgb]{.75,.75,.75}0}\\ {\color[rgb]{.75,.75,.75}0}\\ {\color[rgb]{.75,.75,.75}0}\\ {\color[rgb]{.75,.75,.75}1}\\ {\color[rgb]{.75,.75,.75}0}\end{pmatrix}\quad\rightarrow\quad\begin{pmatrix}{\color[rgb]{0,0,0}1}\\ {\color[rgb]{.75,.75,.75}0}\\ {\color[rgb]{.75,.75,.75}1}\\ {\color[rgb]{.75,.75,.75}1}\\ {\color[rgb]{.75,.75,.75}0}\\ {\color[rgb]{0,0,0}1}\\ {\color[rgb]{.75,.75,.75}0}\\ {\color[rgb]{.75,.75,.75}1}\\ {\color[rgb]{.75,.75,.75}0}\end{pmatrix}\quad\rightarrow\quad\begin{pmatrix}{\color[rgb]{0,0,0}1}\\ {\color[rgb]{.75,.75,.75}0}\\ {\color[rgb]{.75,.75,.75}1}\\ {\color[rgb]{0,0,0}0}\\ {\color[rgb]{.75,.75,.75}0}\\ {\color[rgb]{0,0,0}1}\\ {\color[rgb]{.75,.75,.75}0}\\ {\color[rgb]{.75,.75,.75}1}\\ {\color[rgb]{.75,.75,.75}0}\end{pmatrix}\quad\rightarrow\quad\begin{pmatrix}{\color[rgb]{1,.5,0}1}\\ {\color[rgb]{1,.5,0}0}\\ {\color[rgb]{1,.5,0}0}\\ {\color[rgb]{1,0,0}0}\\ {\color[rgb]{1,0,0}0}\\ {\color[rgb]{1,0,0}1}\\ {\color[rgb]{0,0,1}0}\\ {\color[rgb]{0,0,1}1}\\ {\color[rgb]{0,0,1}0}\end{pmatrix}$
In practice, however, it is often the case that significantly more than three infeasible configurations are required before a new feasible solution is found, as now all feasible solutions exist as local optima, surrounded by high-energy (i.e., penalized by $\lambda$) infeasible solutions. Figure 2.1 illustrates this phenomenon. In Figure 2.1, the tabu-based QUBO heuristic is able to see only 15 feasible solutions, indicated by the line segments, over 1000 iterations. Annealing-based QUBO solvers often have even worse performance, particularly as the temperature decreases and they are forced to accept only moves that result in an improvement.
Figure 2.1: Hamming distance between the solutions observed by three heuristic solvers and the optimal solution for a QAPLIB instance with 15 facilities and locations (225 binary variables in total). The search heuristics used are native space tabu search, a random generator of feasible solutions, and a QUBO space tabu search.
Figure 2.2: Relative optimality gap as a function of the number of iterations for the QAPLIB [27] instances between 12 and 25 facilities and locations (or between 144 and 625 binary variables). Four heuristics, each repeated $30$ times, are tested: a tabu implementation on the QUBO formulation, a tabu implementation on the native formulation, a greedy formulation on the native formulation, and a baseline of generated random feasible solutions.
Solvers attempting to navigate this landscape tend to spend most of the time on infeasible solutions, and are thus able to find only a few feasible solutions. Any notion of performing gradient descent on the feasible space is lost as there is no direct connection between feasible solutions. Figure 2.2 shows that, in comparing optimized solvers in the QUBO space to the generation of random feasible solutions (which is trivial, as feasible solutions are simply permutations), the QUBO heuristics do not improve on the random heuristic. The native-space tabu search heuristic, on the other hand, drastically outperforms the QUBO heuristic and is often able to find the optimal solution.
3 An Improved Local Search Heuristic for Hardware
The QAP problem QAP$(F,D)$ can be formulated as a series of matrix multiplications by representing solutions as permutation matrices. Recall that $P_{n}$ is the set of $n\times n$ permutation matrices. Then, the objective function representing an optimal solution $\pi^{*}$ in the problem (1.1) is equivalent to the formulation
$X^{*}\in\operatorname*{Argmin}_{X\in P_{n}}\left\{\operatorname*{trace}(DXFX^{T})\right\}.$ (3.1)
It is well-known that, for any $n\times n$ matrices $D,F$, and $X$, the following relationship holds:
$\begin{array}[]{rcl}\operatorname*{trace}(FXD^{T}X^{T})&=&\operatorname*{vec}(F^{T}XD)^{T}\operatorname*{vec}(X)\\ &=&\left[(D^{T}\otimes F^{T})\operatorname*{vec}(X)\right]^{T}\operatorname*{vec}(X)\\ &=&\operatorname*{vec}(X)^{T}(D\otimes F)\operatorname*{vec}(X).\end{array}$ (3.2)
Thus, the formulation (3.1) is equivalent to
$X^{*}\in\operatorname*{Argmin}_{X\in P_{n}}\left\{\operatorname*{vec}(X)^{T}(D\otimes F)\operatorname*{vec}(X)\right\},$ (3.3)
and this expression represents the non-penalized part of the binary formulation in the QUBO objective function (2.1).
In this section, we propose a local update mechanism that adheres to the formulation (3.3). We emphasize that working within the constraints
$\left\{\operatorname*{vec}(X)\in\{0,1\}^{n^{2}}:X\in P_{n}\right\}$
removes the need to employ the penalty terms in the QUBO formulation (2.1); from this point forward, we refer to this formulation as the “binary formulation”. For the balance of this paper, we assume that $D$ and $F$ have diagonal elements of zero, as is observed in the vast majority of QAP instances.
3.1 Quadruple Bit-Flip
The main differences in solving a QAP problem in the native space compared to solving it in the QUBO space are that the QUBO solution space is much larger and that it is composed of mostly infeasible solutions. In addition, the basic move on a single bit-flip QUBO solver would force the solver to spend most of its time in infeasible space and impede any gradient decisions based only on feasible solutions. It is possible to avoid the infeasible solutions and regain the performance of a native QAP solver by defining each move as a quadruple bit-flip move within the constrained feasible binary space. The key idea is that flipping a single zero variable in a feasible solution uniquely determines the other three variables that will need to be flipped in order to guarantee another feasible solution.
We illustrate how a quadruple bit-flip is determined with an initially feasible solution with $n=4$ native variables (i.e., 16 variables in the binary space). Suppose that the variable at index seven is flipped from zero to one:
$x=\begin{pmatrix}1\\ 0\\ 0\\ 0\\ 0\\ 0\\ {\color[rgb]{1,0,0}1}\\ 1\\ 0\\ 0\\ 1\\ 0\\ 0\\ 1\\ 0\\ 0\\ \end{pmatrix}\quad\rightarrow\quad\begin{bmatrix}1&0&0&0\\ 0&0&0&1\\ 0&{\color[rgb]{1,0,0}1}&1&0\\ 0&1&0&0\\ \end{bmatrix}.$
Visualizing the binary solution as a matrix immediately exposes the violations of the one-hot encoding constraints (i.e., rows and columns must contain exactly one “1”), and pinpoints the bit-flips that are required to arrive at a new feasible solution:
$x=\begin{pmatrix}1\\ 0\\ 0\\ 0\\ 0\\ 0\\ {\color[rgb]{1,0,0}1}\\ {\color[rgb]{1,0,0}0}\\ 0\\ 0\\ {\color[rgb]{1,0,0}0}\\ {\color[rgb]{1,0,0}1}\\ 0\\ 1\\ 0\\ 0\\ \end{pmatrix}\quad\rightarrow\quad\begin{bmatrix}1&0&0&0\\ 0&0&0&1\\ 0&{\color[rgb]{1,0,0}1}&{\color[rgb]{1,0,0}0}&0\\ 0&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}1}&0\\ \end{bmatrix}.$
For an arbitrary $n$, there are $n^{2}-n=n(n-1)$ variables assigned a value of “0” from which to choose, and a permutation solution in native space has $\binom{n}{2}=n(n-1)/2$ possible local moves. The 2$\times$ disparity comes from the fact that there are two possible choices of “0” elements in the binary solution that result in the identical new feasible solution. Transitioning from one feasible solution to another involves selecting a pair of variables assigned a value of “1”. This choice then uniquely determines the two other variables assigned a value of “0” that need to be flipped. We can systematically perform these operations directly within the feasible binary space. Given two indices $z_{1},z_{2}\in\{0,1,\ldots,n^{2}-1\}$ to flip from “1” to “0”, we identify indices $z_{3},z_{4}\in\{0,1,\ldots,n^{2}-1\}$ to flip from “0” to “1” following the rule summarized in Algorithm 3.1.
Algorithm 3.1   Quadruple Bit-Flips in Feasible Binary Space
1:function Identify_Bit_to_Flip($z_{1},z_{2}$)
2:     $z_{3}=\left\lfloor{z_{2}}/{n}\right\rfloor\cdot n+\left(z_{1}\bmod n\right)$
3:     $z_{4}=\left\lfloor{z_{1}}/{n}\right\rfloor\cdot n+\left(z_{2}\bmod n\right)$
4:     return $z_{3},z_{4}$
5:end function
[…]
$x(z_{1})\leftarrow 0,\ x(z_{2})\leftarrow 0,\ x(z_{3})\leftarrow 1,\ x(z_{4})\leftarrow 1.$
There are $n$ elements assigned a value of “1” in each feasible binary solution $x$ that correspond to $\operatorname*{supp}(x)$. Therefore, the local neighbourhood of possible feasible moves comprises $\binom{n}{2}$ quadruple bit-flips.
3.2 Feasible-Space Updates to the Binary Formulation
In this section, we derive a formula to compute the difference in the objective values between two permutation assignments, expressed through a sequence of inner products. The formula facilitates natural translations to IMC hardware. We use the quadruple bit-flips outlined in Section 3.1 and vectorized operations. Starting from a feasible solution $x\in\{0,1\}^{n^{2}}$, let $y\in\{0,1\}^{n^{2}}$ be a feasible solution derived from $x$ via a quadruple bit-flip. For simplicity, let $Q:=D\otimes F$ in the formulation (3.3). The goal is to show how to compute the gradient, that is, the difference in the objective value for a local update,
$y^{T}Qy-x^{T}Qx,$ (3.4)
using simple matrix–vector products.
Let $z_{1}$ and $z_{2}$ be the indices where “$1$” in $x$ is flipped to “00”, while $z_{3}$ and $z_{4}$ are the indices where “00” in $x$ is flipped to “$1$”, following the procedure outlined in Algorithm 3.1. Since $y$ is the result of the quadruple bit-flip originating from $x$, $\operatorname*{supp}(x)\cap\operatorname*{supp}(y)^{c}=\{z_{1},z_{2}\}$ and $\operatorname*{supp}(y)\cap\operatorname*{supp}(x)^{c}=\{z_{3},z_{4}\}$. We define the binary vectors $\phi_{x},\phi_{y}\in\{0,1\}^{n^{2}}$, where
$(\phi_{x})_{i}=\begin{cases}1&\text{if }i\in\{z_{1},z_{2}\},\\ 0&\text{otherwise,}\end{cases}\quad\text{ and }\quad(\phi_{y})_{i}=\begin{cases}1&\text{if }i\in\{z_{3},z_{4}\},\\ 0&\text{otherwise.}\end{cases}$ (3.5)
In other words, $\phi_{x}=e_{z_{1}}+e_{z_{2}}$ and $\phi_{y}=e_{z_{3}}+e_{z_{4}}$, where $e_{i}$ denotes the $i$-th standard unit vector. The support of the vectors $\phi_{x}$ and $\phi_{y}$ indicates the positions in the incumbent variable assignment where bit flips have occurred. Thus, the two solutions $x$ and $y$ have the following relationship:
$y=x-e_{z_{1}}-e_{z_{2}}+e_{z_{3}}+e_{z_{4}}=x-\phi_{x}+\phi_{y}.$
Then, the difference (3.4) can be computed explicitly as
$\begin{array}[]{ccl}y^{T}Qy-x^{T}Qx&=&\left[Q(z_{3},:)+Q(z_{4},:)\right]y-\left[Q(z_{1},:)+Q(z_{2},:)\right]x\\ &&+\left[Q(:,z_{3})+Q(:,z_{4})\right]^{T}y-\left[Q(:,z_{1})+Q(:,z_{2})\right]^{T}x\\ &&+~{}Q(z_{1},:)\phi_{x}+Q(z_{2},:)\phi_{x}-Q(z_{3},:)\phi_{y}-Q(z_{4},:)\phi_{y}.\end{array}$ (3.6)
The derivation of the equality (3.6) is included in Proposition A.1.
The remainder of this section focuses on reducing the number of algebraic operations outlined in the equality (3.6). We employ symmetrization as a strategy to achieve efficient local updates and present an improved local update that leverages the structure of semi-symmetric instances.
3.2.1 Feasible-Space Updates via Symmetrization
For any $w\in\mathbb{R}^{n}$ and $A\in\mathbb{R}^{n\times n}$, recall the identity
$w^{T}Aw=w^{T}\left(\frac{1}{2}A+\frac{1}{2}A^{T}\right)w.$
Hence, we may symmetrize $Q$ and assume the symmetry of $Q$ without loss of generality. Since we have $Q(:,i)=Q(:,i)^{T}$ holds for all $i$, this symmetrization step significantly reduces the number of inner products in the naïve computation (3.6):
$\begin{array}[]{ccl}y^{T}Qy-x^{T}Qx&=&2\cdot\left[Q(z_{3},:)+Q(z_{4},:)\right]y-2\cdot\left[Q(z_{1},:)+Q(z_{2},:)\right]x\\ &&+~{}Q(z_{1},:)\phi_{x}+Q(z_{2},:)\phi_{x}-Q(z_{3},:)\phi_{y}-Q(z_{4},:)\phi_{y}.\end{array}$ (3.7)
Furthermore, since $Q$ has diagonal elements with a value of “0”, note that
$Q(z_{1},:)\phi_{x}=Q(z_{1},:)e_{z_{1}}+Q(z_{1},:)e_{z_{2}}=Q(z_{1},z_{1})+Q(z_{1},z_{2})=Q(z_{1},z_{2}).$
Similarly, we have
$Q(z_{2},:)\phi_{x}=Q(z_{2},z_{1}),\ Q(z_{3},:)\phi_{y}=Q(z_{3},z_{4}),\ \text{and }\ Q(z_{4},:)\phi_{y}=Q(z_{4},z_{3}).$
Therefore, the last four terms of the right-hand side of the equality (3.7) can be simplified owing to the symmetry of $Q$, resulting in a more compact gradient evaluation:
$\begin{array}[]{ccl}&&y^{T}Qy-x^{T}Qx=2\cdot\left[Q(z_{3},:)+Q(z_{4},:)\right]y+2\cdot\left[-Q(z_{1},:)-Q(z_{2},:)\right]x+2\cdot\left[Q(z_{1},:)\phi_{x}-Q(z_{4},:)\phi_{y}\right].\end{array}$ (3.8)
3.2.2 Feasible-Space Updates for Semi-symmetric Instances
A QAP instance is said to be “symmetric” if both the flow matrix $F$ and the distance matrix $D$ are symmetric, “semi-symmetric” if either $D$ or $F$ is symmetric, and “asymmetric” if neither of the matrices is symmetric. The QAPLIB [27] is a popular benchmarking library widely used by researchers. Out of 136 instances in the QAPLIB, only eight QAP instances formulated by Burkard and Offermann (the “bur” class) are asymmetric. We further simplify the formula (3.8) for symmetric or semi-symmetric instances in Proposition 3.1 below.
Proposition 3.1.
Let $D\in\mathbb{R}^{n\times n}$, $F\in\mathbb{S}^{n}$ and $Q=\frac{1}{2}(D\otimes F)+\frac{1}{2}(D\otimes F)^{T}$. Let $x=\operatorname*{vec}(X)$, where $X\in P_{n}$, and $y=x-e_{z_{1}}-e_{z_{2}}+e_{z_{3}}+e_{z_{4}}$, with $(z_{1},z_{2},z_{3},z_{4})$ being the tuple that realizes the quadruple bit-flip. Then,
$\begin{array}[]{c}y^{T}Qy-x^{T}Qx=2\cdot\left[Q(z_{3},:)y+Q(z_{4},:)y-Q(z_{1},:)x-Q(z_{2},:)x\right].\end{array}$ (3.9)
Proof.
Let $\phi_{x}=e_{z_{1}}+e_{z_{2}}$ and $\phi_{y}=e_{z_{3}}+e_{z_{4}}$. We show that
$Q(z_{1},:)\phi_{x}=Q(z_{4},:)\phi_{y}$ (3.10)
holds, where
$Q=\frac{1}{2}(D\otimes F)+\frac{1}{2}(D\otimes F)^{T}=\left(\frac{1}{2}D+\frac{1}{2}D^{T}\right)\otimes F.$
Let $\hat{D}=\frac{1}{2}D+\frac{1}{2}D^{T}$. Since $F$ is symmetric, each $n\times n$ block submatrix within $\hat{D}\otimes F$ is symmetric:
$\hat{D}_{i,j}F_{k,\ell}=\hat{D}_{i,j}F_{\ell,k},\ \forall i,j,k,\ell\in\{0,\ldots,n-1\}.$
Hence, we have
$Q(in+k,jn+\ell)=Q(in+\ell,jn+k),\ \forall i,j\in\{0,\ldots,n-1\},k,\ell\in\{0,\ldots,n-1\}.$ (3.11)
Let $i=\lfloor z_{1}/n\rfloor$ and $j=\lfloor z_{2}/n\rfloor$. Recall Algorithm 3.1, and note that
$\begin{array}[]{lcl}z_{1}=in+(z_{1}\bmod n),&&z_{2}=jn+(z_{2}\bmod n),\\ z_{3}=jn+(z_{1}\bmod n),&&z_{4}=in+(z_{2}\bmod n).\\ \end{array}$ (3.12)
Thus, $Q(z_{1},z_{2})=Q(z_{4},z_{3})$ holds by the identity (3.11). Now, recall that
$Q(z_{1},:)\phi_{x}=Q(z_{1},z_{2})\ \text{ and }\ Q(z_{4},:)\phi_{y}=Q(z_{4},z_{3}).$
Thus, the equality (3.10) holds and the equality (3.9) follows. ∎
The reduced gradient evaluation in the equality (3.9) is equally applicable to semi-symmetric instances with a symmetric matrix $D$, rather than $F$ being symmetric. The cyclicity of the trace operator provides
$\operatorname*{trace}(FXD^{T}X^{T})=\operatorname*{trace}(D^{T}X^{T}FX).$
Then, by a change of variables, $Z\leftarrow X^{T}$, we obtain
$\operatorname*{trace}(FXD^{T}X^{T})=\operatorname*{trace}(D^{T}ZFZ^{T}).$
Thus, by the equality (3.2), we have
$\operatorname*{vec}(X)^{T}(D\otimes F)\operatorname*{vec}(X)=\operatorname*{vec}(Z)^{T}(F^{T}\otimes D^{T})\operatorname*{vec}(Z)=\operatorname*{vec}(Z)^{T}(F\otimes D)\operatorname*{vec}(Z).$
Algorithm 3.2 shows how to formulate the data matrix $Q$ used in the binary space representation of the QAP (i.e., the QUBO matrix formulation) used throughout this paper.
Algorithm 3.2   QUBO Matrix Formulation
1:function Formulate_QUBO_Matrix($F$leftarrow flow matrix, $D$leftarrow distance matrix)
2:     if $F,D$ are symmetric then $\triangleright$ symmetric instance
3:         $Q=F\otimes D$
4:     else if exactly one of $F,D$ is symmetric then $\triangleright$ semi-symmetric instance
5:         if $F$ is symmetric then $Q=D\otimes F$
6:         else if $D$ is symmetric then $Q=F\otimes D$
7:         end if
8:         $Q=\frac{1}{2}Q+\frac{1}{2}Q^{T}$
9:     else $\ F,D$ are not symmetric $\triangleright$ asymmetric instance
10:         $Q=F\otimes D$
11:         $Q=\frac{1}{2}Q+\frac{1}{2}Q^{T}$
12:     end if
13:     return $Q$
14:end function
4 Full-Neighbourhood Search
In the native QAP space, given a solution represented as a permutation matrix $X\in P_{n}$, we call $Y$ a “neighbour” of $X$ if $Y$ is obtained by swapping two distinct columns of $X$. A “full neighbourhood” of $X$ is the set of all such neighbours of $X$. Hence, the full neighbourhood of $X$ consists of $\binom{n}{2}=\frac{n(n-1)}{2}$ members. In this section, we present an analogous interpretation of the full neighbourhood in the binary space, followed by a vectorized evaluation mechanism for computing the gradients associated with the full neighbourhood in the binary space. Our approach focuses on parallelizing the evaluation of the objective function within a single iteration of the algorithm, rather than utilizing multiple parallel threads to evaluate each individual objective function.
4.1 Definition of “Full Neighbourhood”
Let $x=\operatorname*{vec}(X)$, where $X\in P_{n}$ is a permutation matrix. The full neighbourhood of $x$, denoted by $\mathcal{N}(x)$, is the set of $4$-tuples
$\mathcal{N}(x):=\left\{\left(z_{1},z_{2},\left\lfloor\frac{z_{2}}{n}\right\rfloor\cdot n+\bmod(z_{1},n),\left\lfloor\frac{z_{1}}{n}\right\rfloor\cdot n+\bmod(z_{2},n)\right):z_{1},z_{2}\in\operatorname*{supp}(x),z_{1}<z_{2}\right\}.$ (4.1)
Note that the last two elements in each tuple originate from the quadruple bit-flip defined in Algorithm 3.1. We use the term “neighbour” interchangeably to refer both to an element $(z_{1},z_{2},z_{3},z_{4})\in\mathcal{N}(x)$ and to the vector $y=x-e_{z_{1}}-e_{z_{2}}+e_{z_{3}}+e_{z_{4}}$ when the meaning is clear. Since $\mathcal{N}(x)$ has an equivalent interpretation in the native space $P_{n}$, there are exactly $\binom{n}{2}$ members in $\mathcal{N}(x)$. For a given neighbour $(z_{1},z_{2},z_{3},z_{4})\in\mathcal{N}(x)$, it is important to note that
$\left\lfloor\frac{z_{1}}{n}\right\rfloor=\left\lfloor\frac{z_{4}}{n}\right\rfloor\text{ and }\left\lfloor\frac{z_{2}}{n}\right\rfloor=\left\lfloor\frac{z_{3}}{n}\right\rfloor.$ (4.2)
Hence, in the context of permutation matrices, each neighbour in $\mathcal{N}(x)$ corresponds to swapping the $\left\lfloor\frac{z_{1}}{n}\right\rfloor$-th column with the $\left\lfloor\frac{z_{2}}{n}\right\rfloor$-th column of $X$. Lastly, according to the definition of $\mathcal{N}(x)$ in the set (4.1), the first two elements in each $4$-tuple represent the indices of $x$ assigned a value of “1”. Since $\mathcal{N}(x)$ collects all such pairs $z_{1},z_{2}\in\operatorname*{supp}(x)$, taking their union collectively reconstructs $\operatorname*{supp}(x)$. Recall that the last two elements of each tuple designate the indices of $x$ that need to be flipped from “0” to “1”. Note that these elements are uniquely determined by the first two elements. This observation leads to the following algebraic set relationships:
$\begin{array}[]{lcl}\text{supp}(x)&=&\cup_{(z_{1},z_{2},z_{3},z_{4})\in\mathcal{N}(x)}\{z_{1},z_{2}\},\\ \text{supp}(x)^{c}&=&\cup_{(z_{1},z_{2},z_{3},z_{4})\in\mathcal{N}(x)}\{z_{3},z_{4}\}.\end{array}$ (4.3)
Example 4.1 illustrates the conformations of the full neighbourhood in three different solution spaces.
Example 4.1.
(Equivalence of three configurations) Consider the permutation $\pi=[2,0,1,3,4]\in\Pi_{5}$ and its counterpart $X=$ $\begin{bmatrix}0&1&0&0&0\\ 0&0&1&0&0\\ 1&0&0&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\\ \end{bmatrix}$ realized in $P_{5}$. Then, swapping two elements of $\pi$ is equivalent to swapping two columns of $X$, resulting in the following $10$ permutation matrices accounting for the full neighbourhood of $X$:
$\begin{bmatrix}1&0&0&0&0\\ 0&0&1&0&0\\ 0&1&0&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\\ \end{bmatrix},\begin{bmatrix}0&1&0&0&0\\ 1&0&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\\ \end{bmatrix},\begin{bmatrix}0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 1&0&0&0&0\\ 0&0&0&0&1\\ \end{bmatrix},\begin{bmatrix}0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&0&1\\ 0&0&0&1&0\\ 1&0&0&0&0\\ \end{bmatrix},\begin{bmatrix}0&0&1&0&0\\ 0&1&0&0&0\\ 1&0&0&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\\ \end{bmatrix},$
$\begin{bmatrix}0&0&0&1&0\\ 0&0&1&0&0\\ 1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&0&0&1\\ \end{bmatrix},\begin{bmatrix}0&0&0&0&1\\ 0&0&1&0&0\\ 1&0&0&0&0\\ 0&0&0&1&0\\ 0&1&0&0&0\\ \end{bmatrix},\begin{bmatrix}0&1&0&0&0\\ 0&0&0&1&0\\ 1&0&0&0&0\\ 0&0&1&0&0\\ 0&0&0&0&1\\ \end{bmatrix},\begin{bmatrix}0&1&0&0&0\\ 0&0&0&0&1\\ 1&0&0&0&0\\ 0&0&0&1&0\\ 0&0&1&0&0\\ \end{bmatrix},\begin{bmatrix}0&1&0&0&0\\ 0&0&1&0&0\\ 1&0&0&0&0\\ 0&0&0&0&1\\ 0&0&0&1&0\\ \end{bmatrix}.$
Swapping two elements of $\pi$ has an analogous meaning in the feasible binary space. Let $x=\operatorname*{vec}(X)$. Swapping two elements of $\pi$ at indices 00 and $1$ yields the permutation $[0,2,1,3,4]$. Realizing this swap in the binary space is equivalent to the binary solution $x-e_{2}-e_{5}+e_{7}+e_{0}$. The remaining neighbours in $\mathcal{N}(x)$ can be identified similarly using Table 4.1. Note that $\cup_{(z_{1},z_{2},z_{3},z_{4})\in\mathcal{N}(x)}\{z_{1},z_{2}\}$ collects $\mathrm{supp}(x)=\{2,5,11,18,24\}$ and $\cup_{(z_{1},z_{2},z_{3},z_{4})\in\mathcal{N}(x)}\{z_{3},z_{4}\}$ collects the complement $\mathrm{supp}(x)^{c}=\{0,1,\ldots,24\}\setminus\{2,5,11,18,24\}$.
Swap Indices of $\pi$ Label for Members in $\mathcal{N}(x)$ $z_{1}$ $z_{2}$ $z_{3}$ $z_{4}$
0 $\leftrightarrow$ 1 $N_{0}$ 2 5 7 0
0 $\leftrightarrow$ 2 $N_{1}$ 2 11 12 1
0 $\leftrightarrow$ 3 $N_{2}$ 2 18 17 3
0 $\leftrightarrow$ 4 $N_{3}$ 2 24 22 4
1 $\leftrightarrow$ 2 $N_{4}$ 5 11 10 6
1 $\leftrightarrow$ 3 $N_{5}$ 5 18 15 8
1 $\leftrightarrow$ 4 $N_{6}$ 5 24 20 9
2 $\leftrightarrow$ 3 $N_{7}$ 11 18 16 13
2 $\leftrightarrow$ 4 $N_{8}$ 11 24 21 14
3 $\leftrightarrow$ 4 $N_{9}$ 18 24 23 19
Table 4.1: Full neighbourhood at $x=(0,0,1,0,0,1,0,0,0,0,0,1,0,0,0,0,0,0,1,0,0,0,0,0,1)^{T}$.
4.2 Partial Update of the Full Neighbourhood
Let $\mathcal{N}(x)$ be the full neighbourhood of a feasible solution $x$. We fix a neighbour $(\bar{z}_{1},\bar{z}_{2},z_{2}^{*},z_{1}^{*})\in\mathcal{N}(x)$. Note that by the definition (4.1), the elements in this neighbour are related by the equalities
$z_{1}^{*}=\left\lfloor\frac{\bar{z}_{1}}{n}\right\rfloor\cdot n+\bmod(\bar{z}_{2},n)\ \text{ and }\ z_{2}^{*}=\left\lfloor\frac{\bar{z}_{2}}{n}\right\rfloor\cdot n+\bmod(\bar{z}_{1},n).$ (4.4)
Now, consider the neighbour that results from the quadruple bit-flip via $(\bar{z}_{1},\bar{z}_{2},z_{2}^{*},z_{1}^{*})$,
$y=x-e_{\bar{z}_{1}}-e_{\bar{z}_{2}}+e_{z_{1}^{*}}+e_{z_{2}^{*}},$
and its full neighbourhood $\mathcal{N}(y)$. In this section, we derive $\mathcal{N}(y)$ through a partial update to $\mathcal{N}(x)$.
Recall, from the definition (4.1) and the equalities (4.3), that the first two elements of all members of the full neighbourhoods $\mathcal{N}(x)$ and $\mathcal{N}(y)$ are characterized by all pairs of elements in $\operatorname*{supp}(x)$ and $\operatorname*{supp}(y)$, respectively, while the last two elements are induced by the first two. Since $x$ and $y$ are related by a quadruple bit-flip, we have $|\operatorname*{supp}(x)\cap\operatorname*{supp}(y)|=n-2$. This implies that, among the members in $\mathcal{N}(x)$ and $\mathcal{N}(y)$, those originating from $\operatorname*{supp}(x)\cap\operatorname*{supp}(y)$ are identical; see Figure 4.1(a). Hence, among $\binom{n}{2}$ members in each of $\mathcal{N}(x)$ and $\mathcal{N}(y)$, exactly $\binom{n-2}{2}$ members are shared, that is,
$|\mathcal{N}(x)\cap\mathcal{N}(y)|=\binom{n-2}{2}.$
Consequently, $\binom{n}{2}-\binom{n-2}{2}=2n-3$ members within $\mathcal{N}(x)$ require updates to construct $\mathcal{N}(y)$.
Figure 4.1: (a) Relationship between two feasible binary solutions $x,y\in\{0,1\}^{n^{2}}$. Consequently, there are $\binom{n-2}{2}$ elements in $\mathcal{N}(x)\cap\mathcal{N}(y)$. (b) Disparity between the total number of members in the full neighbourhood and those requiring updates.
Now, we define the set
$\mathcal{U}_{\mathcal{N}(x)}:=\left\{(z_{1},z_{2},z_{3},z_{4})\in\mathcal{N}(x):z_{1}\in\{\bar{z}_{1},\bar{z}_{2}\}\text{ or }z_{2}\in\{\bar{z}_{1},\bar{z}_{2}\}\right\}.$
Note that the $32n-$ 3 members within $\mathcal{N}(x)$ requiring updates are characterized by $\mathcal{U}_{\mathcal{N}(x)}$. Notably, the cardinality of $\mathcal{U}_{\mathcal{N}(x)}$ is linear in $n$, while the cardinality of $\mathcal{N}(x)$ is quadratic in $n$; see Figure 4.1(b).
4.2.1 Updates Realized using the Tuples in $\mathcal{N}(x)$
Let $\dot{\cup}$ denote the disjoint union. We partition $\mathcal{U}_{\mathcal{N}(x)}$ based on the membership of $\bar{z}_{1}$ and $\bar{z}_{2}$ within each tuple:
$\mathcal{U}_{\mathcal{N}(x)}=\mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{1}}\dot{\cup}\ \mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{2}}\dot{\cup}\ \mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{1},\bar{z}_{2}},$
where
$\begin{array}[]{rcl}\mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{1}}&=&\left\{(z_{1},z_{2},z_{3},z_{4})\in\mathcal{N}(x):\bar{z}_{1}\in\{z_{1},z_{2}\},\bar{z}_{2}\not\in\{z_{1},z_{2}\}\right\},\\ \mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{2}}&=&\left\{(z_{1},z_{2},z_{3},z_{4})\in\mathcal{N}(x):\bar{z}_{2}\in\{z_{1},z_{2}\},\bar{z}_{1}\not\in\{z_{1},z_{2}\}\right\},\text{ and }\\ \mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{1},\bar{z}_{2}}&=&\left\{(z_{1},z_{2},z_{3},z_{4})\in\mathcal{N}(x):\{\bar{z}_{1},\bar{z}_{2}\}=\{z_{1},z_{2}\}\right\}.\end{array}$
We note that
$\left|\mathcal{U}_{\mathcal{N}(x)}\right|=\left|\mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{1}}\right|+\left|\mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{2}}\right|+\left|\mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{1},\bar{z}_{2}}\right|=(n-2)+(n-2)+1=2n-3.$
Let $(z_{1},z_{2},z_{3},z_{4})\in\mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{1}}\subseteq\mathcal{N}(x)$ and assume that $z_{1}=\bar{z}_{1}$. We observe the changes made to $(z_{1},z_{2},z_{3},z_{4})$ when $z_{1}$ is updated to $z_{1}^{*}$, while $z_{2}$ remains unchanged. From the definition (4.1), the third element, $z_{3}$, in this tuple is transformed according to
$\left\lfloor\frac{z_{2}}{n}\right\rfloor\cdot n+\bmod(z_{1},n)\rightarrow\left\lfloor\frac{z_{2}}{n}\right\rfloor\cdot n+\bmod(z_{1}^{*},n).$ (4.5)
The fourth element remains unchanged due to the relationship (4.2); hence, we have the equality
$\left\lfloor\frac{z_{1}}{n}\right\rfloor\cdot n+\bmod(z_{2},n)=\left\lfloor\frac{z_{1}^{*}}{n}\right\rfloor\cdot n+\bmod(z_{2},n).$ (4.6)
Thus, it follows from the assignment (4.5) and the equality (4.6) that the neighbour $(z_{1},z_{2},z_{3},z_{4})$ in $\mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{1}}$ is updated as follows:
$(z_{1},z_{2},z_{3},z_{4})\to\left(z_{1}^{*},z_{2},\left\lfloor\frac{z_{2}}{n}\right\rfloor\cdot n+\bmod(z_{1}^{*},n),z_{4}\right).$ (4.7)
Similarly, for $(z_{1},z_{2},z_{3},z_{4})\in\mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{2}}$ with $z_{2}=\bar{z}_{2}$, the neighbour is updated as follows:
$(z_{1},z_{2},z_{3},z_{4})\to\left(z_{1},z_{2}^{*},z_{3},\left\lfloor\frac{z_{1}}{n}\right\rfloor\cdot n+\bmod(z_{2}^{*},n)\right).$ (4.8)
Lastly, for $(z_{1},z_{2},z_{3},z_{4})\in\mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{1},\bar{z}_{2}}$, the neighbour is updated as follows:
$(z_{1},z_{2},z_{3},z_{4})\to\left(z_{1}^{*},z_{2}^{*},\left\lfloor\frac{z_{2}}{n}\right\rfloor\cdot n+\bmod(z_{1}^{*},n),\left\lfloor\frac{z_{1}}{n}\right\rfloor\cdot n+\bmod(z_{2}^{*},n)\right)=(z_{1}^{*},z_{2}^{*},\bar{z}_{2},\bar{z}_{1}).$
Example 4.2.
Recall Example 4.1 and $\mathcal{N}(x)$ in Table 4.1. Let “$N_{7}$” from the full neighbourhood $\mathcal{N}(x)$ corresponding to $(z_{1},z_{2},z_{3},z_{4})=(11,18,16,13)$ be chosen for the quadruple bit-flip. Then, the corresponding new solution is $y=x-e_{11}-e_{18}+e_{16}+e_{13}$. Then, the full neighbourhood at $y$, $\mathcal{N}(y)$, is shown in Table 4.2. Note that $N_{1},N_{2},N_{4},N_{5},N_{7},N_{8},N_{9}$ from $\mathcal{N}(x)$ are updated to produce $\mathcal{N}(y)$ in Table 4.2, while $N_{0},N_{3},N_{6}$ from $\mathcal{N}(x)$ and $\mathcal{N}(y)$ are identical. The changed elements are highlighted in bold in Table 4.2. Of note, exactly two elements in $N_{1},N_{2},N_{4},N_{5},N_{8},N_{9}$ are modified, whereas all elements of $N_{7}$ are affected.
$\mathcal{N}(x)$ $\mathcal{N}(y)$
label $z_{1}$ $z_{2}$ $z_{3}$ $z_{4}$ label $z_{1}$ $z_{2}$ $z_{3}$ $z_{4}$
$N_{0}$ 2 5 7 0 $N_{0}$ 2 5 7 0
$N_{1}$ 2 11 12 1 $N_{1}$ 2 13 12 3
$N_{2}$ 2 18 17 3 $N_{2}$ 2 16 17 1
$N_{3}$ 2 24 22 4 $N_{3}$ 2 24 22 4
$N_{4}$ 5 11 10 6 $N_{4}$ 5 13 10 8
$N_{5}$ 5 18 15 8 $N_{5}$ 5 16 15 6
$N_{6}$ 5 24 20 9 $N_{6}$ 5 24 20 9
$N_{7}$ 11 18 16 13 $N_{7}$ 13 16 18 11
$N_{8}$ 11 24 21 14 $N_{8}$ 13 24 23 14
$N_{9}$ 18 24 23 19 $N_{9}$ 16 24 21 19
Table 4.2: Full neighbourhood $\mathcal{N}(x)$ at the solution $x$ from Example 4.1 and the full neighbourhood $\mathcal{N}(y)$ at $y=(0,0,1,0,0,1,0,0,0,0,0,\textbf{0},0,\textbf{1},0,0,\textbf{1},0,\textbf{0},0,0,0,0,0,1)^{T}$.
4.2.2 Updates Executed Using Matrices
In principle, it is possible to construct $\mathcal{N}(y)$ from scratch by enumerating all neighbours of $x$ directly. However, instead of performing this exhaustive enumeration, we use a more efficient approach by constructing two sparse matrices. These matrices are designed to systematically update the set of neighbours from $\mathcal{N}(x)$ to $\mathcal{N}(y)$. By leveraging these matrices, we can perform targeted neighbour updates, avoiding the need to rebuild the full neighbourhood from the ground up.
Let the members in the set $\mathcal{N}(x)$ be initialized to follow the lexicographic order. The neighbours in Tables 4.1 and 4.2 show examples of the lexicographic ordering. Let the matrices $S_{12},S_{34}\in\{0,1\}^{\binom{n}{2}\times n^{2}}$, where the supports of the $i$-th rows are
$\text{supp}(S_{12}[i,:])=\{z_{1},z_{2}\},\ \ \text{supp}(S_{34}[i,:])=\{z_{3},z_{4}\},\ \ \text{where }(z_{1},z_{2},z_{3},z_{4})\text{ is the {\hbox{i}}-th member of }\mathcal{N}(x).$ (4.9)
Note that $\operatorname*{supp}(x)=\operatorname*{supp}(\mathbf{1}^{T}S_{12})$ and $\{0,\ldots,n^{2}-1\}\setminus\operatorname*{supp}(x)=\operatorname*{supp}(\mathbf{1}^{T}S_{34})$.
In order to detect the rows of $S_{12}$ and $S_{34}$ that correspond to the neighbours in $\mathcal{U}_{\mathcal{N}(x)}$, let $b\in\{0,1,2\}^{n^{2}}$, where
$b_{i}=\begin{cases}1&\text{if }i=\bar{z}_{1},\\ 2&\text{if }i=\bar{z}_{2},\\ 0&\text{otherwise. }\end{cases}$ (4.10)
We then use $S_{12}b$ to detect the rows of $S_{12}$ that contain the members of $\mathcal{U}_{\mathcal{N}(x)}$. Consider the following sets:
$\mathcal{R}_{1}:=\{i:(S_{12}b)_{i}=1\},\ \mathcal{R}_{2}:=\{i:(S_{12}b)_{i}=2\},\ \text{ and }\ \mathcal{R}_{3}:=\{i:(S_{12}b)_{i}=3\}.$ (4.11)
The set $\mathcal{R}_{1}$ contains the row indices of $S_{12}$ and $S_{34}$ that correspond to the neighbours $\mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{1}}$. Similarly, $\mathcal{R}_{2}$ and $\mathcal{R}_{3}$ contain the row indices of $S_{12}$ and $S_{34}$ corresponding to the neighbours $\mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{2}}$ and $\mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{1},\bar{z}_{2}}$, respectively. Of note, $|\mathcal{R}_{1}|=|\mathcal{R}_{2}|=|\mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{1}}|=|\mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{2}}|=n-2$ and $|\mathcal{R}_{3}|=|\mathcal{U}_{\mathcal{N}(x)}^{\bar{z}_{1},\bar{z}_{2}}|=1$.
Now we update $S_{12}$ to represent the first two elements in each $4$-tuple of $\mathcal{N}(y)$ as follows:
  1. 1.
    For all $i\in\mathcal{R}_{1}$, let $S_{12}(i,\bar{z}_{1})=0$ and $S_{12}(i,z_{1}^{*})=1$;
  2. 2.
    For all $i\in\mathcal{R}_{2}$, let $S_{12}(i,\bar{z}_{2})=0$ and $S_{12}(i,z_{2}^{*})=1$; and
  3. 3.
    For $i\in\mathcal{R}_{3}$, let $S_{12}(i,[\bar{z}_{1},\bar{z}_{2}])=0$ and $S_{12}(i,[z_{1}^{*},z_{2}^{*}])=1$.
Note that this update involves changing $22n-$ 2 elements of $S_{12}$ from 0 to 1, and another $22n-$ 2 changes from 1 to 0. See Figure 4.2 (a), (b), and (c) for an illustration of the transition of this update for an instance of size $n=12$. The update procedure for $S_{12}$ is outlined in Algorithm 4.1.
Figure 4.2: The transition of the full-neighbourhood lookup matrices $S_{12},S_{34}$ at the permutation assignments $\pi_{1}=[3,\textbf{1},2,0,4,7,6,\textbf{5},8,9,10,11]\in\Pi_{12}$ and $\pi_{2}=[3,\textbf{5},2,0,4,7,6,\textbf{1},8,9,10,11]\in\Pi_{12}$. The assignment $\pi_{2}$ is obtained after swapping the elements of $\pi_{1}$ at indices $1$ and $7$. (a) and (d) show the full neighbourhood of $\pi_{1}$; (b) and (e) show the full neighbourhood of $\pi_{2}$. (c) highlights the difference between (a) and (b), and (f) highlights the between (d) and (e). Among $266=n(n-1)/$ 2 members in the full neighbourhood of $\pi_{1}$, $321=2n-$ 3 members need updates, resulting in $44=4(n-1)$ element changes in both $S_{12}$ and $S_{34}$.
Algorithm 4.1   Partial Neighbourhood Update for $S_{12}$
1:function Partial_Neighbourhood_Update_S12($z_{1},z_{2},z_{1}^{*},z_{2}^{*}$, $S_{12}$, $\mathcal{R}_{1},\mathcal{R}_{2},\mathcal{R}_{3}$)
2:     $S_{12}(i,z_{1})=0,S_{12}(i,z_{1}^{*})=1$, $i\in\mathcal{R}_{1}$
3:     $S_{12}(i,z_{2})=0,S_{12}(i,z_{2}^{*})=1$, $i\in\mathcal{R}_{2}$
4:     $S_{12}(i,[z_{1},z_{2}])=0,S_{12}(i,[z_{1}^{*},z_{2}^{*}])=1$, $i\in\mathcal{R}_{3}$
5:     return $S_{12}$
6:end function
Recall that a neighbour has an analogous meaning in the space of permutation matrices. Once a neighbour is chosen from $\mathcal{N}(\operatorname*{vec}(X))$, it analogously means that two columns of $X\in P_{n}$ are swapped while the other $n-2$ columns remain invariant. The indices of these invariant column $X$ are characterized by the set
$\mathcal{C}_{0}=\{0,1,2,...,n-1\}\setminus\left\{\left\lfloor\frac{\bar{z}_{1}}{n}\right\rfloor,\left\lfloor\frac{\bar{z}_{2}}{n}\right\rfloor\right\}.$
Given a set $\mathcal{D}$ of numbers, we define the function $\texttt{List}_{\texttt{sorted}}(\mathcal{D})$ by
[…]
Let $c=\texttt{List}_{\texttt{sorted}}(\mathcal{C}_{0})\in\mathbb{R}^{n-2}$. To update the rows in $\mathcal{R}_{1}$ of $S_{34}$, recall the assignments (4.5) and (4.7). The vector
$c_{1}=n\cdot c+(\bar{z}_{1}\bmod n)$
indicates the columns of $S_{34}$ that need to be updated from a value of “1” to “0”. Similarly, the column indices of $S_{34}$ associated with $\mathcal{R}_{2}$ that require updating from “1” to “0” are identified using
$c_{2}=n\cdot c+(\bar{z}_{2}\bmod n).$
As for the coordinates of $S_{34}$ that require updating from “0” to “1”, recall (4.4) and note that
$(\bar{z}_{2}\bmod n)-(z^{*}_{1}\bmod n)+(\bar{z}_{1}\bmod n)=(\bar{z}_{1}\bmod n).$
Thus, the elements of $S_{34}$ with columns corresponding to $n\cdot c+(z^{*}_{1}\bmod n)$ and rows associated with $\mathcal{R}_{1}$ can be identified using $c_{2}$ and $\mathcal{R}_{1}$. Similarly, $c_{1}$ and $\mathcal{R}_{2}$ can be used to identify the coordinates of $S_{34}$ that require updating from “0” to “1”. As with $S_{12}$, updating $S_{34}$ involves changing $44n-$ 4 elements: $22n-$ 2 elements changing from “0” to “1”, and another $22n-$ 2 changing from “1” to “0”.
Let $r_{1}=\texttt{List}_{\texttt{sorted}}(\mathcal{R}_{1})$ and $r_{2}=\texttt{List}_{\texttt{sorted}}(\mathcal{R}_{2})$. Flipping the values of $S_{34}$ at the following coordinates provides the required update:
$\{(r_{1}(i),c_{1}(i))\}_{i\in\{0,1,\ldots,n-3\}},$$\quad\{(r_{1}(i),c_{2}(i))\}_{i\in\{0,1,\ldots,n-3\}},$ (4.12a)
$\{(r_{2}(i),c_{1}(i))\}_{i\in\{0,1,\ldots,n-3\}},$$\quad\{(r_{2}(i),c_{2}(i))\}_{i\in\{0,1,\ldots,n-3\}},$ (4.12b)
$\{(i,\bar{z}_{1}),(i,\bar{z}_{2})\}_{i\in\mathcal{R}_{3}},$$\quad\{(i,z_{1}^{*}),(i,z_{2}^{*})\}_{i\in\mathcal{R}_{3}}.$ (4.12c)
The procedure for updating $S_{34}$ is outlined in Algorithm 4.2, and the transition of $S_{34}$ is illustrated in Figure 4.2 (d), (e), and (f) for an instance of size $n=12$. The full-neighbourhood update executed using arrays is summarized in Algorithm 4.3. The notation $\leftrightarrow$” in Algorithm 4.2 stands for the swap operator. We note that following the update governed by Algorithm 4.3 preserves the lexicographic ordering of the subsequent full neighbourhoods.
Algorithm 4.2   Partial Neighbourhood Update for $S_{34}$
1:function Partial_Neighbourhood_Update_S34($\bar{z}_{1},\bar{z}_{2},z_{1}^{*},z_{2}^{*}$, $S_{34}$, $r_{1},r_{2},\mathcal{R}_{3}$)
2:     $\mathcal{C}_{0}\leftarrow\{0,1,2,...,n-1\}\setminus\left\{\left\lfloor\frac{\bar{z}_{1}}{n}\right\rfloor,\left\lfloor\frac{\bar{z}_{2}}{n}\right\rfloor\right\}$
3:     $c\leftarrow\texttt{List}_{\texttt{sorted}}(\mathcal{C}_{0})$
4:     $c_{1}\leftarrow n\cdot c+(\bar{z}_{2}\bmod n)$
5:     $c_{2}\leftarrow n\cdot c+(\bar{z}_{1}\bmod n)$
6:     $S_{34}(r_{1}[i],c_{1}[i])\leftrightarrow S_{34}(r_{1}[i],c_{2}[i])$,  $S_{34}(r_{1}[i],c_{2}[i])\leftrightarrow S_{34}(r_{1}[i],c_{1}[i])$, $i=0,1,\ldots,n-3$triangleright (4.12a)
7:     $S_{34}(r_{2}[i],c_{1}[i])\leftrightarrow S_{34}(r_{2}[i],c_{2}[i])$,  $S_{34}(r_{2}[i],c_{2}[i])\leftrightarrow S_{34}(r_{2}[i],c_{1}[i])$, $i=0,1,\ldots,n-3$triangleright (4.12b)
8:     $S_{34}(i,[\bar{z}_{1},\bar{z}_{2}])=1,S_{34}(i,[z_{1}^{*},z_{2}^{*}])=0$, $i\in\mathcal{R}_{3}$triangleright (4.12c)
9:     return $S_{34}$
10:end function
Algorithm 4.3   Update of Full Neighbourhood
1:function Update_Full_Neighbourhood($S_{12}$, $S_{34}$, $z_{1},z_{2}$)
2:     $z_{2}^{*}\leftarrow\left\lfloor\frac{z_{2}}{n}\right\rfloor\cdot n+\bmod(z_{1},n)$, $z_{1}^{*}\leftarrow\left\lfloor\frac{z_{1}}{n}\right\rfloor\cdot n+\bmod(z_{2},n)$
3:     $b\leftarrow e_{z_{1}}+2e_{z_{2}}$triangleright The vector (4.10)
4:     $\mathcal{R}_{1}\leftarrow\{i:(S_{12}b)_{i}=1\},\ \mathcal{R}_{2}\leftarrow\{i:(S_{12}b)_{i}=2\},\ \mathcal{R}_{3}\leftarrow\{i:(S_{12}b)_{i}=3\}$triangleright The sets (4.11)
5:     $S_{12}\leftarrow$ Partial_Neighbourhood_Update_S12($z_{1},z_{2},z_{1}^{*},z_{2}^{*}$, $S_{12}$, $\mathcal{R}_{1},\mathcal{R}_{2},\mathcal{R}_{3}$) $\triangleright$ Algorithm 4.1
6:     $r_{1}\leftarrow\texttt{List}_{\texttt{sorted}}(\mathcal{R}_{1})$, $r_{2}\leftarrow\texttt{List}_{\texttt{sorted}}(\mathcal{R}_{2})$
7:     $S_{34}\leftarrow$ Partial_Neighbourhood_Update_S34($z_{1},z_{2},z_{1}^{*},z_{2}^{*}$, $S_{34}$, $r_{1},r_{2},\mathcal{R}_{3}$) $\triangleright$ Algorithm 4.2
8:     return $S_{12},S_{34}$
9:end function
4.3 Vectorized Gradient Evaluation of the Full Neighbourhood for
Symmetric and Semi-symmetric Instances
Given a feasible $x$ and its neighbour $y$, recall, from Proposition 3.1, that semi-symmetric instances allow for the simplification
$\langle y,Qy\rangle-\langle x,Qx\rangle=2\cdot\left[Q(z_{3},:)y+Q(z_{4},:)y-Q(z_{1},:)x-Q(z_{2},:)x\right].$
In this section, we focus on symmetric and semi-symmetric instances. Let $y^{i}$ be the feasible solution realized by the $i$-th element of $\mathcal{N}(x)$. We aim to compute the vector $\Delta\in\mathbb{R}^{\binom{n}{2}}$, where the $i$-th component of $\Delta$ is
[…]
Each $\Delta_{i}$ represents the gradient of the $i$-th neighbour of $\mathcal{N}(x)$. In our discussion, we omit the scalar multiplier “$2$” from Proposition 3.1 since it does not affect the relative ordering of the elements in $\Delta$.
Given a vector $x\in\{0,1\}^{n^{2}}$, we use a matrix $P_{\operatorname*{supp}(x)}:=I_{n^{2}}(\operatorname*{supp}(x),:)\in\{0,1\}^{|\operatorname*{supp}(x)|\times n^{2}}$ that is capable of pulling the elements of $x$ associated with the support $\operatorname*{supp}(x)$. Let $K=I_{n}\otimes\mathbf{1}_{n}\in\{0,1\}^{n^{2}\times n}$. Then, for any vector $x$ with $\operatorname*{supp}(x)=\{i_{1},i_{2},\ldots,i_{n}\}$, $KP_{\operatorname*{supp}(x)}w$ copies the elements of $w$ corresponding to $\operatorname*{supp}(x)$ exactly $n$ times, that is,
$K(P_{\operatorname*{supp}(x)}w)=(\underbrace{w_{i_{1}},w_{i_{1}},\ldots,w_{i_{1}}}_{n\text{ times}},\underbrace{w_{i_{2}},w_{i_{2}},\ldots,w_{i_{2}}}_{n\text{ times}},\ldots,\underbrace{w_{i_{n}},w_{i_{n}},\ldots,w_{i_{n}}}_{n\text{ times}})^{T}.$ (4.13)
We decompose $\Delta_{i}$ into two parts:
$\Delta_{i}=\tilde{\Delta}_{i}+\mathcal{E}_{i},$ (4.14)
where
$\begin{array}[]{rcl}\tilde{\Delta}_{i}&=&Q(z_{3},:)x+Q(z_{4},:)x-Q(z_{1},:)x-Q(z_{2},:)x,\text{ and }\\ \mathcal{E}_{i}&=&Q(z_{3},:)(y^{i}-x)+Q(z_{4},:)(y^{i}-x).\end{array}$ (4.15)
Propositions 4.3 and 4.4 show how the elements $\tilde{\Delta}_{i}$ and $\mathcal{E}_{i}$ associated with each neighbour in $\mathcal{N}(x)$ can be computed using simple matrix–vector operations.
Proposition 4.3.
Let $\mathcal{N}(x)$ be the full neighbourhood of $x=\operatorname*{vec}(X)$, where $X\in P_{n}$. Let $S_{34}$ be the array satisfying the equalities (4.9). Then, the $i$-th element of the vector
$S_{34}g$ (4.16)
yields $\tilde{\Delta}_{i}$ defined in the equality (4.15), where
$\begin{array}[]{rcl}g=-K(P_{\operatorname*{supp}(x)}((Qx)\circ x))+(Qx)\circ(\mathbf{1}_{n^{2}}-x).\\ \end{array}$ (4.17)
Proof.
Let $\bar{x}$ be the negation of the current variable assignment $x$, that is,
$\bar{x}=\mathbf{1}_{n^{2}}-x.$
We fix a row index $i$ of $S_{34}$. Observe that
$e_{i}^{T}S_{34}\left((Qx)\circ\bar{x}\right)=(e_{z_{3}}^{T}+e_{z_{4}}^{T})\left((Qx)\circ\bar{x}\right)=Q(z_{3},:)x+Q(z_{4},:)x.$
Let $\operatorname*{supp}(x)=\{i_{1},i_{2},\ldots,i_{n}\}$. Then,
$K(P_{\operatorname*{supp}(x)}((Qx)\circ x))=(\underbrace{(Qx)_{i_{1}},(Qx)_{i_{1}},\cdots,(Qx)_{i_{1}}}_{n\text{ times}},\ldots,\underbrace{(Qx)_{i_{n}},(Qx)_{i_{n}},\ldots,(Qx)_{i_{n}}}_{n\text{ times}})^{T}.$
Let $\{z_{1},z_{2}\}=\operatorname*{supp}(e_{i}^{T}S_{12})$ be the corresponding counterpart of $\{z_{3},z_{4}\}$. The relationship (4.2) implies
$e_{i}^{T}S_{34}\left(K(P_{\operatorname*{supp}(x)}((Qx)\circ x))\right)=Q(z_{1},:)x+Q(z_{2},:)x.$
Then, with the vector
$g=-K(P_{\operatorname*{supp}(x)}((Qx)\circ x))+(Qx)\circ\bar{x},$
one can verify that $\tilde{\Delta}=S_{34}g$. ∎
We note that the matrix $I_{n}\otimes\mathbf{1}_{n}^{T}$ in the proof above can be used instead of the projection $P_{\operatorname*{supp}(x)}$. Thus, a fixed matrix $(I_{n}\otimes\mathbf{1}_{n})(I_{n}\otimes\mathbf{1}_{n}^{T})=I_{n}\otimes(\mathbf{1}_{n}\mathbf{1}_{n}^{T})$ can be used instead of $KP_{\operatorname*{supp}(x)}$. We illustrate the computation of the vector $g$ from the equality (4.17) in Figure 4.3.
Figure 4.3: Diagram of the computation of the equality (4.17) at variable assignment $x=\operatorname*{vec}\left(\begin{bmatrix}e_{2}&e_{0}&e_{1}&e_{3}&e_{4}\end{bmatrix}\right)$.
Proposition 4.4.
Let $\mathcal{N}(x)$ be the full neighbourhood of $x=\operatorname*{vec}(X)$, where $X\in P_{n}$. Let $S_{34}$ be the array satisfying the equality (4.9). Then, the $i$-th element of the vector
$\begin{array}[]{rcl}\mathcal{E}=\left((S_{34}Q)\circ S_{34}\right)\mathbf{1}_{n^{2}}\end{array}$ (4.18)
yields the equality (4.15).
Proof.
For a fixed $i$-th neighbour $(z_{1},z_{2},z_{3},z_{4})\in\mathcal{N}(x)$, we observe that
$\begin{array}[]{rcl}\mathcal{E}_{i}&=&Q(z_{3},:)(y^{i}-x)+Q(z_{4},:)(y^{i}-x)\\ &=&Q(z_{3},:)(\phi_{y}-\phi_{x})+Q(z_{4},:)(\phi_{y}-\phi_{x})\\ &=&(Q(z_{3},:)+Q(z_{4},:))\phi_{y}-(Q(z_{3},:)+Q(z_{4},:))\phi_{x},\end{array}$ (4.19)
where $\phi_{x}$ and $\phi_{y}$ are as defined in the equalities (3.5).
We first consider the terms $(Q(z_{3},:)+Q(z_{4},:))\phi_{y}$ in the equality (4.19). For each $i$, observe that
$e_{i}^{T}S_{34}Qe_{j}=Q(z_{3},j)+Q(z_{4},j),\text{ where }\{z_{3},z_{4}\}=\operatorname*{supp}(e_{i}^{T}S_{34}).$
Then, the $i$-th element of the column sum of $((S_{34}Q)\circ S_{34})$ is
$e_{i}^{T}\left((S_{34}Q)\circ S_{34}\right)\mathbf{1}_{n^{2}}=Q(z_{3},z_{3})+Q(z_{3},z_{4})+Q(z_{4},z_{3})+Q(z_{4},z_{4})\\ =Q(z_{3},z_{4})+Q(z_{4},z_{3}).$
Thus, $e_{i}^{T}\left((S_{34}Q)\circ S_{34}\right)\mathbf{1}_{n^{2}}=(Q(z_{3},:)+Q(z_{4},:))\phi_{y}$.
We now consider $(Q(z_{3},:)+Q(z_{4},:))\phi_{x}$ in the equality (4.19). We aim to show that $Q(z_{3},:)\phi_{x}=0$ and $Q(z_{4},:)\phi_{x}=0$. Recall the algebraic relationship (3.12) among $z_{1},z_{2},z_{3}$, and $z_{4}$. For simplicity, we let
$\begin{array}[]{lcl}z_{1}=kn+t_{1},&&z_{2}=jn+t_{2},\\ z_{3}=jn+t_{1},&&z_{4}=kn+t_{2}.\\ \end{array}$
Then, we may represent $e_{z_{1}}=e_{kn+t_{1}}\in\{0,1\}^{n^{2}}$ using the Kronecker product
$e_{z_{1}}=e_{k}\otimes e_{t_{1}},\text{ where }e_{k},e_{t_{1}}\in\{0,1\}^{n}.$
Similarly, with $e_{k},e_{j},e_{t_{1}},e_{t_{2}}\in\{0,1\}^{n}$, we have
$\begin{array}[]{lcl}e_{z_{2}}=e_{j}\otimes e_{t_{2}},\ e_{z_{3}}=e_{j}\otimes e_{t_{1}},\text{ and }e_{z_{4}}=e_{k}\otimes e_{t_{2}}.\end{array}$
Note that $Q(z_{3},:)\phi_{x}=e_{z_{3}}^{T}Qe_{z_{1}}+e_{z_{3}}^{T}Qe_{z_{2}}$. We let $Q=D\otimes F$ without loss of generality. Now observe the expansion
$\begin{array}[]{rcl}e_{z_{3}}^{T}Qe_{z_{1}}&=&(e_{j}\otimes e_{t_{1}})^{T}(D\otimes F)(e_{k}\otimes e_{t_{1}})\\ &=&(e_{j}\otimes e_{t_{1}})^{T}(De_{k}\otimes Fe_{t_{1}})\\ &=&(e_{j}^{T}De_{k})\otimes(e_{t_{1}}^{T}Fe_{t_{1}})\\ &=&0,\end{array}$
where the last equality holds, since $F$ is assumed have diagonal elements of zero. Similarly, since $D$ is also assumed to have diagonal elements of zero, we have
$\begin{array}[]{rcl}e_{z_{3}}^{T}Qe_{z_{2}}&=&(e_{j}\otimes e_{t_{1}})^{T}(D\otimes F)(e_{j}\otimes e_{t_{2}})\\ &=&(e_{j}\otimes e_{t_{1}})^{T}(De_{j}\otimes Fe_{t_{2}})\\ &=&(e_{j}^{T}De_{j})\otimes(e_{t_{1}}^{T}Fe_{t_{2}})\\ &=&0.\end{array}$
Using a similar expansion, one can obtain the equality $Q(z_{4},:)\phi_{x}=0$. Therefore, the element $e_{i}^{T}\mathcal{E}$ is equal to $Q(z_{3},:)(y^{i}-x)+Q(z_{4},:)(y^{i}-x)$. ∎
The steps for computing $\Delta$ using Propositions 4.3 and 4.4 is summarized in Algorithm 4.4. We view $\tilde{\Delta}$ as an approximate gradient and $\Delta$ as the exact gradient. Algorithm 4.4 is illustrated in Example 4.5, with a problem size of $n=5$.
Algorithm 4.4   Full-Neighbourhood Evaluation
1:function Evaluate_Full_Neighbourhood($x$, $S_{12},S_{34}$)
2:     $g\leftarrow K(P_{\operatorname*{supp}(x)}((Qx)\circ x))-(Qx)\circ(\mathbf{1}_{n^{2}}-x)$triangleright The equality (4.17)
3:     $\Delta\leftarrow S_{34}g$triangleright The approximate gradient (4.16)
4:     if exact gradient is desired then
5:         $\Delta\leftarrow\Delta+\left((S_{34}Q)\circ(S_{34})\right)\mathbf{1}_{n^{2}}$triangleright The exact gradient using the error corrector (4.18)
6:     end if
7:     return $\Delta$
8:end function
Example 4.5.
(Illustration of Algorithm 4.4) We continue with the permutation assignment $\pi=[2,0,1,3,4]$ from Example 4.1. Consider the following distance and flow matrices:
$D=\begin{bmatrix}0&1&0&1&4\\ 1&0&4&2&0\\ 0&0&0&3&5\\ 5&3&1&0&1\\ 4&5&0&3&0\\ \end{bmatrix}\in\mathbb{R}^{5\times 5},\quad F=\begin{bmatrix}0&2&1&3&4\\ 2&0&3&2&2\\ 1&3&0&7&0\\ 3&2&7&0&3\\ 4&2&0&3&0\\ \end{bmatrix}\in\mathbb{S}^{5}.$
Let $Q=\frac{1}{2}(D\otimes F)+\frac{1}{2}(D\otimes F)^{T}$ be formed via Algorithm 3.2. Then
$K(P_{\operatorname*{supp}(x)}((Qx)\circ x))=(I_{5}\otimes\mathbf{1}_{5})\begin{pmatrix}11.5&15&13&20.5&21\end{pmatrix}^{T}\in\mathbb{R}^{25},$
and
$\begin{array}[]{ccrrrrrrrrrrrrr}(Qx)\circ(\mathbf{1}_{25}-x)&=&(16&13&0&12&8.5&0&13&23.5&18.5&11.5&10&0\ldots\\ &&16&7.5&14&13.5&13.5&8.5&0&14&9&21&24&33&0)^{T}\in\mathbb{R}^{25}.\end{array}$
Hence, the vector $g$ in the equality (4.17) is
$\begin{array}[]{ccrrrrrrrrrrrrrr}g&=&(4.5&1.5&-11.5&0.5&-3&-15&-2&8.5&3.5&-3.5&-3&-13\ldots\\ &&3&-5.5&1&-7&-7&-12&-20.5&-6.5&-12&0&3&12&-21)^{T}\in\mathbb{R}^{25}.\end{array}$
The computations of the vector (4.16), representing the approximate gradient $S_{34}g$, and the equality (4.18), representing the error correction vector $\mathcal{E}$, are
[…]
The exact gradient is computed by $\Delta=\tilde{\Delta}+\mathcal{E}=\begin{pmatrix}15&4.5&9.5&0&3&-3.5&4.5&-4.5&11&17.5\end{pmatrix}^{T}$.
4.4 Parallel Full-Neighbourhood Search Algorithm
We summarize all of the routines introduced in Section 3, Section 4.2, and Section 4.3 in Algorithm 4.5.
Algorithm 4.5   Parallel Full-Neighbourhood Local Search Algorithm
1:function Full_Neighbourhood_Local_Search($F\leftarrow$ flow matrix, $D\leftarrow$ distance matrix, $X\leftarrow$ permutation matrix, $i_{\text{max}}$leftarrow maximum number of iterations)
2:     $Q\leftarrow\texttt{Formulate\_QUBO\_Matrix}(F,D)$triangleright Algorithm 3.2
3:     $x\leftarrow\operatorname*{vec}(X)$
4:     $(f_{\text{best}},x_{\text{best}})\leftarrow(\langle x,Qx\rangle,x)$
5:     Initialize $S_{12},S_{34}$triangleright The equality (4.9)
6:     $i\leftarrow 0$
7:     while $i<$ $i_{\text{max}}$  do
8:         $\Delta\leftarrow$ Evaluate_Full_Neighbourhood($x$, $S_{12}$, $S_{34}$) $\triangleright$ Algorithm 4.4
9:         $(z_{1},z_{2},z_{3},z_{4})\leftarrow$ Choose_Move($S_{34},\Delta$) $\triangleright$ A heuristic of choice
10:         $x(z_{1})\leftarrow 0,\ x(z_{2})\leftarrow 0,\ x(z_{3})\leftarrow 1,\ x(z_{4})\leftarrow 1$triangleright A quadruple bit-flip
11:         if $\langle x,Qx\rangle<f_{\text{best}}$ then
12:              $(f_{\text{best}},x_{\text{best}})\leftarrow(\langle x,Qx\rangle,x)$
13:         end if
14:         $S_{12},S_{34}\leftarrow$ Update_Full_Neighbourhood($S_{12},S_{34},z_{1},z_{2}$) $\triangleright$ Algorithm 4.3
15:         $i\leftarrow i+1$
16:     end while
17:     return $(f_{\text{best}},x_{\text{best}})$
18:end function
We make a few remarks pertaining to Algorithm 4.5. The explicit heuristic is not specified the function Choose_Move (line 9 of Algorithm 4.5). This flexibility is a strength of our approach and allows for the choosing of the most appropriate local search heuristic. This adaptability ensures that the approach can be easily integrated with a wide range of heuristics available in the literature. The array $S_{34}$ is passed as an input to Choose_Move because, once a value $\Delta_{i}$ provides the gradient, the corresponding neighbour must be identified for the quadruple bit-flip. The computational cost associated with the error correction term $\left((S_{34}Q)\circ S_{34})\right)\mathbf{1}_{n^{2}}$ is significant in the function Evaluate_Full_Neighbourhood (line 8 of Algorithm 4.5) primarily due to the presence of the matrix–matrix product $S_{34}Q$. As such, we may consider omitting the error correction routine and rely on the vector (4.16) which is the approximation. There are two rationales behind this. Each $\mathcal{E}_{i}$ involves two elements of $Q$. Meanwhile, given $y^{i}$ derived from the $i$-th member $(z_{1},z_{2},z_{3},z_{4})$ of $\mathcal{N}(x)$, the exact gradient is computed via
$\Delta_{i}=Q(z_{3},:)y^{i}+Q(z_{4},:)y^{i}-Q(z_{1},:)x-Q(z_{2},:)x.$
Note that $\Delta_{i}$ involves $4n$ elements of $Q$. Therefore, by excluding $\mathcal{E}_{i}$, only two out of $4n$ terms are miscalculated. As the problem order $n$ increases, the impact on these two error terms diminishes. At the same time, if the algorithm always chooses the move that leads to the best immediate reduction in the value of the objective function, it many become trapped in a local minimum. Hence, a heuristic mechanism is needed to help the algorithm escape from local minima. Moreover, selecting a good candidate from a full-neighbourhood search requires relative ordering of values in $\Delta$, rather than the values themselves. We perform benchmarking with respect to the impact of omitting the error correction steps in Section 5.2 below.
5 Numerical Experiments
We used instances from the QAPLIB [27] of various sizes ranging from $n=12$ to $n=100$ to evaluate the performance of Algorithm 4.5. All instances from the QAPLIB follow a size-based naming convention, “identifier_$n$_version”, where $n$ indicates the problem size. The relative optimality gap was chosen as the performance metric. Let
[…]
We let $f^{*}$ be the best known objective value obtained from the QAPLIB. We define the relative optimality gap as a function of the maximum allowed iterations:
$\operatorname*{\text{rel\_gap}}(i_{\text{max}}):=\frac{f_{\text{best}}(i_{\text{max}})-f^{*}}{f^{*}}.$
We let $\mathcal{T}$ represent the set of trials, or repeats, that is, the number of times an instance is called by Algorithm 4.5. The best optimality gap is then quantifiable by the minimum of the relative optimality gap:
$\operatorname*{\text{rel\_gap}_{\text{min}}}(i_{\text{max}}):=\min_{t\in\mathcal{T}}\left\{{\operatorname*{\text{rel\_gap}}}^{t}(i_{\text{max}})\right\},\text{ for }i_{\text{max}}\in\mathbb{N}.$
We evaluate our approach in two settings of gradient computations. Let $x$ be a feasible binary solution with its neighbour $y=x-e_{z_{1}}-e_{z_{2}}+e_{z_{3}}+e_{z_{4}}$, where $(z_{1},z_{2},z_{3},z_{4})\in\mathcal{N}(x)$. One approach uses the exact gradient $\Delta$ using the equality (4.14), while the other employs the approximation $\tilde{\Delta}$ (see Proposition 4.3). In the context of Section 4.3, we evaluate gradients in $\tilde{\Delta}+\mathcal{E}$ or $\tilde{\Delta}$. We use the symbol $\Theta$ to mean the following:
$\Theta=\begin{cases}\tilde{\Delta}+\mathcal{E}&\text{if the exact gradient is used,}\\ \tilde{\Delta}&\text{if the approximate gradient is used.}\end{cases}$
5.1 Heuristics
We evaluate four heuristics that can be employed in the Choose_Move function (line 9 in Algorithm 4.5). The first heuristic, that we label “Greedy”, is the greedy approach. Given a set of candidates $\mathcal{N}(x)$ and their corresponding value of $\Theta$, we select the one that results in the smallest value, that is,
[…]
The second heuristic is a pseudo-greedy approach, that we label “Top10”. This heuristic randomly chooses one neighbour from the set of neighbours that yields the top 10 best elements in $\Theta$. Let $\widehat{\Theta}=\texttt{List}_{\texttt{sorted}}(\Theta)$, with $\Theta$ computed using $\mathcal{N}(x)$. Let $\pi\in\Pi\left(\binom{n}{2}\right)$ be the permutation that yields the ordered vector $\widehat{\Theta}=\Theta(\pi)$. We then select a neighbour at random:
$N^{*}=\texttt{Top10}(x)\in\left\{\pi^{-1}(0),\pi^{-1}(1),\ldots,\pi^{-1}(9)\right\}.$ (5.2)
Third, we use a random walk heuristic inspired by the well-known WalkSAT heuristic [28] in the area of satisfiability problems. Given a probability $p\in(0,1)$, this heuristic chooses a neighbour based on the following rule:
$N^{*}=\texttt{WalkQAP}(x):=\begin{cases}\texttt{Top10}(x)&\text{with the probability $p$},\\ \text{randomly selected neighbour from $\mathcal{N}(x)$}&\text{with the probability $1-p$.}\end{cases}$ (5.3)
Finally, we use the tabu search heuristic [31], employing the label “Tabu”. In brief, a tabu search constructs a so-called tabu list, which records the feasible points visited up to a given limit. Once $\widehat{\Theta}=\texttt{List}_{\texttt{sorted}}(\Theta)$ is computed, this heuristic sequentially iterates over the neighbours within the full neighbourhood. If a neighbour $i$ results in the value of $f(x)+\widehat{\Theta}_{i}$ being better than the best objective value found so far, we select that neighbour for the next iteration. Otherwise, we choose the incumbent neighbour, provided it is not in the tabu list. Once a neighbour is selected, it is appended to the tabu list, and if the tabu list is at its size limit, the oldest item in the list is removed.
5.1.1 Typical Behaviour of the Greedy Search Heuristic
Given a feasible binary solution $x$, the greedy heuristic iteratively selects the neighbour within $\mathcal{N}(x)$ that yields the maximum immediate reduction in the objective function. While this strategy ensures the best possible improvement at each step within the incumbent full neighbourhood, if often leads to an oscillatory pattern, where the search alternates between two permutation assignments. As a result, the algorithm fails to make further progress towards a global optimal solution. We observe this pattern for five of the selected QAPLIB instances in Figure 5.1. We make observations for both settings of $\Theta$, namely, $\Delta$ and $\tilde{\Delta}$.
Figure 5.1: Convergence of the Greedy heuristic on five QAPLIB instances over 50 iterations of full-neighbourhood evaluations.
From the seventh iteration of QAPLIB instance “chr12a”, when the exact gradient is used, the greedy search alternates between the permutation assignments $[5,6,11,1,9,0,7,10,2,8,4,3]\in\Pi_{12}$ and $[5,6,11,1,9,0,7,8,2,10,4,3]\in\Pi_{12}$, yielding objective values of 14,464 and 14,430, respectively. Similarly, when the approximate gradient is used, the greedy search alternate between $[2,6,9,5,11,8,10,3,0,1,4,7]\in\Pi_{12}$ and $[2,6,9,5,11,8,10,3,1,0,4,7]\in\Pi_{12}$, starting from the first iteration, yielding objective values $50266$ and $48284$, respectively. Such behaviour is a commonly observed pattern and it indicates that the greedy search heuristic struggles to escape local minima. Consequently, we do not consider greedy search for further analysis.
5.2 Comparison of Exact and Approximate Gradients
In this section, we discuss the results of our experiments using the Top10 heuristic (5.2) for two settings of $\Theta$. The number of repeats was set to $100$, that is, $|\mathcal{T}|=100$. The Intel Xeon Gold 6548Y+ processor with 128 cores and 504 GB of memory was used for our experiments. Our aim is to assess the impact of omitting the error correction term $\mathcal{E}$ presented in Proposition 4.4. The primary motivation for omitting this step stems from its computational cost. Computing $\mathcal{E}$ involves a matrix–matrix multiplication $S_{34}Q$, which requires more-substantial processing resources than a matrix–vector multiplication. These operations significantly increase the overall computation time, especially as the problem size grows. The second reason is that when it comes to choosing a neighbour in a given full neighbourhood, the relative ordering of the values in $\Theta$ is important, not the values themselves. For example, given $\Theta^{1}=(-70,-40,-100,3,5)$ and $\Theta^{2}=(-80,-50,-90,3,5)$ representing two distinct gradients, we select the neighbour associated with the smallest value, which is the third element of each vector. Finally, IMC hardware devices typically have limited bit-precision capabilities. As a result, quantization of the nonzero elements in $Q$ becomes necessary. This introduces another potential source of error due to the loss of precision. Consequently, it is crucial to design algorithms that are resilient to multiple sources of error. Despite these challenges, incorporating the error correction step may still be feasible in certain scenarios, particularly when matrix–matrix multiplications can be efficiently performed.
Figure 5.2: (a) Range of relative optimality gaps between approximate and exact gradient computations across $100$ trials with $10^{5}$ total iterations. (b) Best optimality gaps for approximate and exact gradient computations across $10^{4}$ and $10^{5}$ total iterations.
Figure 5.2(a) shows a comparison of the relative optimality gaps using the approximate gradient $\tilde{\Delta}$ and exact gradient $\Delta$ observed over $10^{5}$ iterations, that is, $\left\{{\operatorname*{\text{rel\_gap}}}^{t}(10^{5}):t\in\mathcal{T}\right\}$. Figure 5.2(b) shows the best observed optimality gaps, $\operatorname*{\text{rel\_gap}_{\text{min}}}(10^{4})$ and $\operatorname*{\text{rel\_gap}_{\text{min}}}(10^{5})$, for both $\tilde{\Delta}$ and $\Delta$. Both approaches yield optimality gaps that are smaller than $1\%$, demonstrating their effectiveness in reaching optimal or near-optimal solutions. This observation supports the use of the approximate gradient as a viable alternative to the exact gradient. However, when comparing the relative runtime of computing the exact versus approximate gradients, the advantage of using the approximate gradient becomes evident. Figure 5.3 shows the relative runtime, computed as the ratio of the time taken for $\Delta$ divided by the time taken for $\tilde{\Delta}$ for each instance, with the instances sorted by size. The plot shows that using the exact gradient requires significantly more computational resources and takes considerably longer than the approximate gradient. Notably, the relative time increase becomes more pronounced as the instance size grows.
Figure 5.3: Ratio of the average runtime using the exact gradient versus the approximate gradient across $100$ trials.
We assess the impact of omitting the error terms $\mathcal{E}_{i}$ in the gradient computation as the problem size $n$ increases. For each $i$-th neighbour in $\mathcal{N}(x)$ at a given feasible binary solution $x$, recall that each $\Delta_{i}$ involves $4n$ elements of $Q$, whereas $\mathcal{E}_{i}$ is determined by two elements of $Q$ regardless of the problem size. To ensure a comprehensive evaluation, three classes from the QAPLIB that encompass a wide range of problem sizes are selected. The magnitude of the error terms at a given point is quantified using the following relative error metric. Given a $i$-th neighbour $(z_{1},z_{2},z_{3},z_{4})$ in $\mathcal{N}(x)$, we first decompose $\tilde{\Delta}=S_{34}g$ into the positive and negative components, that is, $\tilde{\Delta}=\tilde{\Delta}_{i}^{+}-\tilde{\Delta}_{i}^{-}$, where $\tilde{\Delta}_{i}^{+}=S_{34}((Qx)\circ(\mathbf{1}-x)$ and $\tilde{\Delta}_{i}^{-}=S_{34}\left(K(P_{\operatorname*{supp}(x)}((Qx)\circ x))\right)$. We then measure the relative error:
[…]
Figure 5.4 shows rel_error over the first $1000$ iterations for instances formulated by Skorin-Kapov (“sko”), Li and Pardalos (“lipa”), and Taillard (“tai”). We observe a consistent trend across all three classes, indicating that the contribution of $\mathcal{E}_{i}$ relative to $\Delta_{i}$ diminishes as the problem size increases, hence, the impact of omitting the error terms becomes negligible in large-scale instances.
Figure 5.4: Relative errors, rel_error, defined by the equality (5.4), observed from sko, lipa, and tai instance classes from the QAPLIB for the first $1000$ iterations of Algorithm 4.5.
5.3 Comparison of Different Heuristics
In this section, we assess the effectiveness of the approximate gradient computation using the heuristics Top10, WalkQAP, and Tabu. The Intel Xeon E7-8890 v3 running at 2.50 GHz with 144 cores and 2.95 TB of memory was used for the experiment. For the WalkQAP heuristic, we set the probability parameter $p=0.95$. For the Tabu heuristic, we set the maximum length of the tabu list to $20$, with the oldest entry being removed when a new neighbour is added whenever the list exceeds this limit. Figure 5.5 presents the relative optimality gaps $\operatorname*{\text{rel\_gap}_{\text{min}}}(10^{5})$ for selected instances from the QAPLIB. The results indicate that the Top10 and WalkQAP heuristics achieve better performance compared to the Tabu heuristic.
[…]
5.4 Computation Breakdown of the Parallel Search and Its Applications to IMC Hardware
In this section, we present a computational breakdown of our Python-based CPU implementation of Algorithm 4.5 to identify the computationally intensive routines by varying the problem size from $n=12$ to $n=100$. Specifically, we compare the computation load distribution of the subroutines in Algorithm 4.5 relative to the total runtime. The instance names are given in each pie chart in Figure 5.6, followed by the average runtime for $10$ repetitions in terms of the wall-clock time. The routines for computing $g$ and $\tilde{\Delta}$ (line 14 in Algorithm 4.5) are labelled “$g$ computation” and “$\tilde{\Delta}$ computation”, respectively. The update of the full neighbourhood (line 8 in Algorithm 4.5) is labelled “Neighbourhood $\mathcal{N}$ update”, and all other routines are collectively labelled “Other”. The comparison is Figure 5.6 was generated using a MacBook Pro equipped with an Apple M3 chip and 16 GB of memory. The Python profiler named “cProfile” was used to provide the performance breakdown.
Figure 5.6: Computation load distribution on six QAPLIB instances. The runtimes over $10^{4}$ iterations appear below the instance names in the pie charts.
In Figure 5.6, we observe a clear trend in which the fraction of the computational time required for calculating $g$ and $\tilde{\Delta}$ increases as the problem size grows. This increase in time is due primarily to the fact that these operations involve matrix–vector multiplications. As the problem size increases, the matrix–vector multiplications become more demanding due to the quadratic growth of the binary formulation. Given that these steps are computational bottlenecks, IMC hardware presents a promising opportunity for considerable improvements in terms of computation time. Unlike conventional processors, which suffer from frequent memory access overhead, IMC architectures perform computations directly within memory. As shown in Figure 4.3, our approach is well-suited for this paradigm. Furthermore, the most demanding computation shown in Figure 5.6 is the $\tilde{\Delta}$ computation, which consists of a simple matrix–vector multiplication. Additionally, the energy efficiency of IMC hardware is a notable by-product, as it reduces the energy overhead of conventional digital computers.
6 Conclusions
To sustain the computational performance growth rate in the post-Moore’s law era, new technologies such as IMC have emerged. Mapping the native formulation of binary constrained optimization problems onto the unconstrained binary formulation incurs a massive overhead and remains the primary obstacle limiting the performance of special-purpose hardware. Hence, pairing special-purpose hardware with algorithms that solve optimization problems in their original, native formulation has the potential to surpass current state-of-the-art classical algorithms implemented on a CPU. In this study, we designed an IMC-compatible algorithm to tackle the QAP, which is a well-known binary constrained optimization problem. The one-shot gradient evaluation framework for the QAP can be integrated with state-of-the-art local-search heuristic algorithms. We emphasized the importance of preserving the native embedding of the problem, focusing exclusively on the native feasible space. A key strength of the proposed methodology is its ability to evaluate all gradients within a single iteration of the algorithm. Moreover, we provided a blueprint for implementing application-driven analog hardware specifically designed for the full-neighbourhood local-search heuristic for the QAP. Another advantage of the approach is its flexibility, allowing it to seamlessly incorporate various heuristics available in the literature without requiring significant modifications. Our experimental results demonstrated competitive performance, achieving close proximity to optimality even when the gradients are evaluated approximately. We showed that using the approximate gradient allows for a fast implementation, without sacrificing the quality of the results, as measured through the optimality gap. Furthermore, we highlighted the substantial speedup potential offered by the use of IMC technology. Our results revealed an opportunity for acceleration when leveraging IMC technology, presenting a compelling case for its use across a wide range of industrially relevant problems.
Acknowledgements
We thank our editor, Marko Bucyk, for his careful review and editing of the manuscript. This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) through Air Force Research Laboratory Agreement No. FA8650-23-3-7313. The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.
Appendix A Supplementary Proof
Proposition A.1.
Let $x$ be a feasible solution in the binary space and let $y=x-e_{z_{1}}-e_{z_{2}}+e_{z_{3}}+e_{z_{4}}$ be a feasible solution formed by the quadruple bit-flip on $x$. Then, the difference in the objective function value between $x$ and $y$ is
$\begin{array}[]{ccl}y^{T}Qy-x^{T}Qx&=&\left[Q(z_{3},:)+Q(z_{4},:)\right]y+\left[-Q(z_{1},:)-Q(z_{2},:)\right]x\\ &&+\left[Q(:,z_{3})+Q(:,z_{4})\right]^{T}y+\left[-Q(:,z_{1})-Q(:,z_{2})\right]^{T}x\\ &&+~{}Q(z_{1},:)\phi_{x}+Q(z_{2},:)\phi_{x}-Q(z_{3},:)\phi_{y}-Q(z_{4},:)\phi_{y}.\end{array}$
Proof.
Let $x$ be a feasible solution in the binary space, and let $y$ be a feasible solution after a local update, that is, $y=x-e_{z_{1}}-e_{z_{2}}+e_{z_{3}}+e_{z_{4}}$. Let $\phi_{x}$ and $\phi_{y}$ be as defined in the equalities (3.5). We decompose the indices as follows:
$\mathcal{I}:=(\operatorname*{supp}(x)\cup\operatorname*{supp}(y))\setminus\{z_{1},z_{2},z_{3},z_{4}\}.$
Let $e_{\mathcal{I}}=\sum_{i\in\mathcal{I}}e_{i}$. Then,
$x=\phi_{x}+e_{\mathcal{I}}=e_{z_{1}}+e_{z_{2}}+e_{\mathcal{I}}\ \text{ and }\ y=\phi_{y}+e_{\mathcal{I}}=e_{z_{3}}+e_{z_{4}}+e_{\mathcal{I}}.$
We then have
$\begin{array}[]{ccl}y^{T}Qy-x^{T}Qx&=&y^{T}Q\left(e_{z_{3}}+e_{z_{4}}+e_{\mathcal{I}}\right)-x^{T}Q\left(e_{z_{1}}+e_{z_{2}}+e_{\mathcal{I}}\right)\\ &=&y^{T}Qe_{z_{3}}+y^{T}Qe_{z_{4}}+y^{T}Qe_{\mathcal{I}}-x^{T}Qe_{z_{1}}-x^{T}Qe_{z_{2}}-x^{T}Qe_{\mathcal{I}}\\ &=&y^{T}Qe_{z_{3}}+y^{T}Qe_{z_{4}}+e_{z_{3}}^{T}Qe_{\mathcal{I}}+e_{z_{4}}^{T}Qe_{\mathcal{I}}+e_{\mathcal{I}}^{T}Qe_{\mathcal{I}}\\ &&-~{}x^{T}Qe_{z_{1}}-x^{T}Qe_{z_{2}}-e_{z_{1}}^{T}Qe_{\mathcal{I}}-e_{z_{2}}^{T}Qe_{\mathcal{I}}-e_{\mathcal{I}}^{T}Qe_{\mathcal{I}}.\end{array}$
After cancelling out the term $e_{\mathcal{I}}^{T}Qe_{\mathcal{I}}$, we may write $y^{T}Qy-x^{T}Qx=D_{y}-D_{x}$, where
$\begin{array}[]{rcl}D_{y}&:=&y^{T}Qe_{z_{3}}+y^{T}Qe_{z_{4}}+e_{z_{3}}^{T}Qe_{\mathcal{I}}+e_{z_{4}}^{T}Qe_{\mathcal{I}},\text{ and }\\ D_{x}&:=&x^{T}Qe_{z_{1}}+x^{T}Qe_{z_{2}}+e_{z_{1}}^{T}Qe_{\mathcal{I}}+e_{z_{2}}^{T}Qe_{\mathcal{I}}.\end{array}$
Let
$\begin{array}[]{rcl}O_{y}&:=&e_{z_{3}}^{T}Q\phi_{y}-e_{z_{3}}^{T}Q\phi_{y}+e_{z_{4}}^{T}Q\phi_{y}-e_{z_{4}}^{T}Q\phi_{y},\text{ and }\\ O_{x}&:=&e_{z_{1}}^{T}Q\phi_{x}-e_{z_{1}}^{T}Q\phi_{x}+e_{z_{2}}^{T}Q\phi_{x}-e_{z_{2}}^{T}Q\phi_{x}.\end{array}$
Note that $O_{x}=O_{y}=0$. Expanding $D_{y}$ and $D_{x}$ using $O_{y}$ and $O_{x}$, we obtain the expressions
$\begin{array}[]{rcl}D_{y}&=&D_{y}+O_{y}\\ &=&y^{T}Qe_{z_{3}}+y^{T}Qe_{z_{4}}+e_{z_{3}}^{T}Qy+e_{z_{4}}^{T}Qy-e_{z_{3}}^{T}Q\phi_{y}-e_{z_{4}}^{T}Q\phi_{y},\end{array}$
and
$\begin{array}[]{rcl}D_{x}&=&D_{x}+O_{x}\\ &=&x^{T}Qe_{z_{1}}+x^{T}Qe_{z_{2}}+e_{z_{1}}^{T}Qx+e_{z_{2}}^{T}Qx-e_{z_{1}}^{T}Q\phi_{x}-e_{z_{2}}^{T}Q\phi_{x}.\end{array}$
Thus, $y^{T}Qy-x^{T}Qx=D_{y}-D_{x}$ results in the equality in the statement. ∎